### v1

# Umberto Collodel Central Bank of Malta

Abstract
This paper develops a novel method to simulate financial market reactions to European Central Bank (ECB) press conferences using a Large Language Model (LLM). We create a behavioral, agent-based simulation of 30 synthetic traders, each with distinct risk preferences, cognitive biases, and interpretive styles. These agents forecast Euro interest rate swap levels at 3-month, 2-year, and 10-year maturities, with the variation across forecasts serving as a measure of market uncertainty or disagreement. We evaluate three prompting strategies, naive, few-shot (enriched with historical data), and an advanced iterative “LLM-as-a-Judge” framework, to assess the effect of prompt design on predictive performance. Even the naive approach generates a strong correlation (~0.5) between synthetic disagreement and actual market outcomes, particularly for longer-term maturities. The LLM-as-a-Judge framework further improves accuracy at the first iteration. These results demonstrate that LLM-driven simulations can capture interpretive uncertainty beyond traditional measures, providing central banks with a practical tool to anticipate market reactions, refine communication strategies, and enhance financial stability.
Keywords: large language models, monetary policy surprises, monetary policy communication
JEL Classification: E52, E58, C63
1 Introduction
“It is hard to imagine that we could train a computer to read Federal Reserve transcripts the way we do. […] We thoroughly expect to be made largely redundant by computers eventually, but perhaps not for a few years to come.” Romer and Romer, (2023)
Recent advancements in large language models (LLMs) have opened up a plethora of possibilities for economic and financial research. Due to the way these models are trained, on vast and diverse textual data, and their ability to generate contextually rich and realistic responses, LLMs can now be viewed as implicit computational representations of human behavior. Much like the economic construct of homo economicus, these models can be endowed with information, placed in hypothetical scenarios, and observed for their responses (Horton,, 2023). Notwithstanding their growing capabilities, the application of LLMs in finance has so far been relatively limited. [^1] A notable exception is, for example, Lopez-Lira and Tang, (2023). Yet, the financial domain may be one of the most promising areas for their use. Unlike traditional economic models that assume rational agents and efficient markets, real-world financial markets are often driven by sentiment, narrative, and behavioral biases, factors that are difficult to capture with purely mathematical models, especially in high-frequency settings. [^2] Agent-based models (ABMs) have emerged as a flexible framework to incorporate such heterogeneity (Farmer and Foley,, 2009). In the context of monetary policy, ABMs have been used to simulate expectation formation, market volatility, and policy transmission (Delli Gatti et al.,, 2011). However, most ABMs rely on stylized behavioral rules imposed by the researcher (Horton,, 2023).
A key area where LLMs can offer significant value is in modeling reactions to central bank communication, a cornerstone of modern monetary policy. It is well-established that the effectiveness of monetary policy hinges not just on the decisions themselves, but on the language used and its interpretation by economic agents. Press conferences, speeches, and official statements play a pivotal role in shaping market expectations and, consequently, asset prices (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019). In previous research, we moved beyond the general direction of the market response to highlight the crucial role of response heterogeneity; in the Euro area, high disagreement following a statement transmits to financial markets through a separate channel compared to the standard monetary policy surprise channel, and can dampen the intended impact of policy tightening by lowering the term premium of the yield curve (Collodel and Kunzmann,, 2025).333(Bauer et al.,, 2021) find similar results for the US. Capturing this nuanced, belief-driven heterogeneity is a challenge for traditional models but is a natural fit for the capabilities of LLMs.
This paper introduces a novel approach to modeling market reactions to ECB communication by leveraging the capabilities of LLMs. We develop a behavioral, agent-based simulation of 30 synthetic traders through Gemini LLM, one of Google’s leading AI products. These agents respond to the texts of ECB press conferences by forecasting Euro interest rate swap levels across three key maturities, short (3m), medium (2y), and long-term (10y). The main idea is that the model can capture a range of possible interpretations of these conferences. Thus, the variation in the resulting forecasts, measured by their standard deviation, serves as a proxy for market uncertainty or disagreement. By correlating this synthetic measure with actual market outcomes, we study the ability of LLMs to replicate the heterogeneity of investors’ beliefs.
To further explore how the design of prompting influences this correlation, we implement three distinct prompting strategies. The first is a naive or ”zero-shot” prompt (Radford et al.,, 2019), which serves as the baseline: each agent is assigned a simple, pre-defined level of risk aversion, a set of cognitive biases, and a unique interpretive style. The second is a ”few-shot” prompt (Brown et al.,, 2020), which enriches the agent’s context by including the past three ECB press conferences along with the corresponding changes in market-based disagreement resulting from them. Finally, we introduce an LLM-as-a-Judge prompt, our most advanced strategy (Zheng et al.,, 2023). In this iterative setup, a separate LLM, the Judge, acts as a critical evaluator, systematically reviewing the analyst prompt and its resulting in-sample correlation score. The Judge then rewrites the prompt to improve alignment with market reactions. This revised prompt is fed back into the Analyst LLM, forming a feedback loop that continues over different iterations. The different prompt versions are subsequently evaluated out-of-sample to assess their robustness and decrease concerns relative to model over-fitting.
We find that LLM-based agents can effectively replicate the complex and varied reactions of investors to ECB press conferences, achieving a remarkable correlation of around 0.5 between the synthetic disagreement measure and actual market outcomes for medium- and long-term maturities, a substantial result given the inherent complexity and noise of financial markets. Contrary to expectations, a more detailed few-shot prompt does not improve this correlation; however, the LLM-as-a-Judge framework, significantly enhances predictive accuracy upon its initial application, with performance degrading thereafter. This breakthrough offers policymakers a powerful new tool to preemptively evaluate and refine central bank communications, identifying language that may trigger market disagreement and volatility well before public release. By moving beyond traditional sentiment analysis, our approach provides a nuanced understanding of how monetary policy signals are perceived and transmitted through financial markets. Integrating insights from macroeconomic communication, behavioral finance, and artificial intelligence, this framework presents a transformative lens for both scholars and decision-makers to better anticipate and manage the financial market impact of monetary policy announcements.
2 Literature Review
The paper contributes to three intersecting literatures: monetary policy communication, behavioral expectations, and the emerging use of large language models (LLMs) in economic analysis.
Our first contribution is to the literature on monetary policy communication, particularly the strand that identifies policy shocks using high-frequency movements in asset prices within narrow windows around policy announcements. Seminal studies for the United States show that Federal Reserve statements influence asset prices not only through policy rate changes but also through the qualitative content of the accompanying communication (Gurkaynak et al.,, 2005; Gertler and Karadi,, 2015; Hansen and McMahon,, 2016; Bauer and Swanson,, 2022). In these studies, changes in risk-free market rates measured before and after announcements serve as proxies for monetary policy shocks. In the euro area, similar evidence shows that ECB press conferences contain significant informational content beyond the policy decision itself (Altavilla et al.,, 2019). Recent work by Bauer et al., (2021) further distinguishes between the direction of policy surprises and the uncertainty surrounding the policy stance, showing that elevated uncertainty depresses equity prices, raises sovereign spreads, and increases exchange rate volatility, a pattern we also find for the Euro area (Collodel and Kunzmann,, 2025). Furthermore, Tillmann, (2020) finds that hawkish surprises are less effective in periods of heightened short- and medium-term uncertainty, as investors shift toward longer maturities in search of safety, thereby flattening the yield curve. While this literature has convincingly demonstrated that communication shocks move markets, it remains silent on how they do so. By relying on market-price changes as statistical proxies, the existing approach leaves the interpretive process of agents as a “black box.” Our work opens that box: we construct an agent-based model in which heterogeneous participants directly process the natural language of policy announcements, allowing us to trace how communication shapes expectations, generates uncertainty, and propagates into financial market dynamics.
The second body of work we engage with concerns behavioral expectations. The rational expectations hypothesis, foundational to much of modern macro-finance, has been increasingly challenged by empirical evidence documenting systematic biases and bounded rationality among investors (Barberis and Thaler,, 2003; Bordalo et al.,, 2018). Standard representative-agent models struggle to reconcile such behavioral deviations with observed phenomena like excess volatility and momentum. Agent-based models (ABMs) have emerged as a flexible alternative (Farmer and Foley,, 2009), allowing for heterogeneous, adaptive agents whose interactions generate emergent macroeconomic patterns. Within monetary policy research, ABMs have been used to model expectation formation and policy transmission in bottom-up settings (Delli Gatti et al.,, 2011). Yet two key limitations remain. First, behavioral rules are typically imposed ex ante as stylized heuristics (Horton,, 2023), rather than learned or emergent. Second, conventional ABMs cannot interpret qualitative information directly, despite the fact that monetary policy signals are largely communicated in natural language. This omission bypasses an essential channel of transmission. Our framework addresses both limitations by enabling agents to process and act upon policy-relevant textual information.
Finally, our study connects to a rapidly developing literature applying LLMs to economic analysis. By processing vast amounts of unstructured text, ranging from macroeconomic reports and corporate disclosures to central bank statements, LLMs extend the reach of empirical economics beyond numerical datasets. In forecasting applications, LLMs can be used directly (Castro and Leibovici,, 2023; Hansen et al.,, 2024; Carriero et al.,, 2025) or indirectly via sentiment extraction, where the derived sentiment serves then as an input to other models. [^4] For example, recent work at the ECB demonstrates that ChatGPT-generated sentiment from PMI commentary improves GDP nowcasts relative to traditional benchmarks (de Bondt and Sun,, 2025). In financial markets, fine-tuned models such as FinBERT or OPT have achieved substantially higher predictive accuracy and trading performance than dictionary-based approaches (BIS,, 2024; Kirtac and Germano,, 2024).
A second application is the construction of LLM-powered simulated economic agents, “homo silicus” (Horton,, 2023)—who are endowed with preferences, information, and decision rules, and can interact in artificial market environments. Frameworks like EconAgent (Zhang et al.,, 2023) and ASFM (Li et al.,, 2024) produce heterogeneous, adaptive behavior that is richer and more realistic than the stylized rules of conventional ABMs, while also capturing the interpretation of complex qualitative signals.
A third and especially relevant strand applies LLMs to the analysis of central bank communication. This includes large-scale classification of monetary policy documents by topic, stance, and sentiment, as in the IMF’s recent work on over 75,000 documents from 169 central banks worldwide (Gelos et al.,, 2025), as well as the development of domain-specific models such as CB-LMs, which improve detection of nuanced policy stances (BIS,, 2025), and CentralBankRoBERTa, which classifies macroeconomic agents and sentiment toward them (Gabriel et al.,, 2023). Other advances, like the Bundesbank’s MILA system, employ role-based prompt engineering to analyze ECB statements sentence-by-sentence, providing interpretable assessments of tone (Bundesbank,, 2024). These innovations have revealed systematic shifts in policy stances and measurably improved the prediction of policy decisions (DIW,, 2025).
Our approach delivers a methodological advance with clear policy relevance. We provide the first micro-founded framework in which heterogeneous, boundedly rational agents interpret central bank communication in real time and translate qualitative policy signals into market-relevant actions. This design closes a long-standing gap between the empirical identification of communication shocks and the theoretical modeling of expectation formation. By combining sentiment extraction, agent-based modeling, and LLM-driven language interpretation in a unified environment, we move beyond reduced-form event studies and open the “black box” of how policy language is processed by market participants. The framework is adaptable across institutional settings, enabling comparative analyses of communication strategies by central banks worldwide. This makes it directly relevant to current debates on transparency, forward guidance, and the role of central bank credibility in shaping financial conditions in both advanced and emerging market economies.
3 Methodology
Our methodology proceeds in three main steps: (i) construction of a structured database of ECB communication events and corresponding market reactions; (ii) development of LLM-based synthetic agents embedded in an agent-based simulation framework; and (iii) evaluation of the correspondence between synthetic measures of belief heterogeneity and actual market outcomes. We focus on European Central Bank (ECB) press conferences, which constitute the primary vehicle for conveying monetary policy decisions in the Euro area. Following the literature on high-frequency identification of monetary policy shocks (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019), we define event dates as the days of regularly scheduled ECB Governing Council monetary policy meetings, limiting the sample to those that include both a written policy statement and a subsequent press conference by the ECB President. For each event, we collect the full transcript of the press conference as published on the ECB website as well as the associated OIS post-conference volatility for the short (3-months), medium (2-years), and long (10-years) term. [^5] Following the monetary policy communication literature, we use OIS rates as they reflect pure interest rate expectations without the credit risk component present in government bond yields (Gurkaynak et al.,, 2005). Our final sample comprises 283 ECB conferences between 1998-06-09 and 2025-04-17.
Figure 1 illustrates market-based disagreement following ECB Governing Council meetings across all three maturities. The data reveal pronounced spikes during periods of heightened policy uncertainty, notably the 2008-2009 financial crisis, the European sovereign debt crisis, and the recent tightening cycle initiated in 2022. Volatility exhibits a clear maturity structure, with consistently higher levels at longer tenors reflecting greater uncertainty about the medium- and long-term policy trajectory compared to near-term rate expectations.
Figure 1: Post ECB Governing Council OIS volatility
Note: Percentage points. Volatility is measured as the average min-max range of OIS rates in the 3 days following ECB press conferences. Quarterly averages displayed. For more details, see Collodel and Kunzmann, (2025).
3.1 Agents
We construct a population of 30 synthetic agents, each instantiated through Gemini 2.5-Flash model, one of Google’s most recent large language model. [^6] We use 30 agents to balance computational and cost efficiency with sufficient heterogeneity to capture cross-sectional variation. The agents are designed to interpret the natural language of ECB press conferences and produce quantitative forecasts of Euro IRS rates at the three maturities. The simulation proceeds as follows. Each agent is endowed with:
  1. 1.
    Risk aversion parameter – fixed at the start of the simulation, drawn from a pre-defined distribution to induce heterogeneity in risk preferences.
  2. 2.
    Cognitive bias profile – encompassing tendencies such as optimism, pessimism, anchoring, or overreaction, assigned to influence interpretation.
  3. 3.
    Interpretive style – a qualitative descriptor shaping the reading of ECB communication (e.g., focusing on inflation outlook, growth projections, or geopolitical risks).
These attributes are embedded in the prompt given to the LLM for each agent, ensuring that the same press conference produces different interpretations across agents. [^7] We set the temperature parameter equal to 1. Temperature is a parameter that controls the randomness of a language model’s output. Lower values (e.g., 0.2) make the model more deterministic and focused, while higher values (e.g., 1.0 or above) increase creativity and variability by making the model more likely to sample less probable words.
3.2 Prompting Strategies
We test three prompting strategies of increasing complexity:
Zero-Shot Prompt:
The baseline approach, where the agent receives only the ECB press conference transcript along with its fixed attributes.
Few-Shot Prompt:
Augments the baseline by including the transcripts of the three most recent ECB press conferences and the associated historical changes in market-based disagreement, providing a short memory of past communication–market interactions. [^8] Big LLMs are notoriously difficult to fine-tune because of computational and cost-related issues. Moreover, we do not have enough observations for this use case. In these instances, a ”few-shot” technique has been shown to improve over a ”zero-shot” benchmark (Brown et al.,, 2020).
LLM-as-a-Judge Prompt:
Introduces a second LLM, the “Judge,” which reviews each agent’s initial prompt and in-sample performance (correlation with market disagreement) and iteratively rewrites the prompt to improve alignment. The improved prompt is fed back into the “Analyst” LLM, forming an adaptive feedback loop. The process continues over multiple iterations before all the prompts provided by the analyst are tested out-of-sample. [^9] The split train-test is 75/25 in chronological order. [^10] Following the literature, we use for the ”Judge” a more potent model compared to the ”Analyst”, namely Gemini 2.5-Pro [^11] See Algorithm 1 for the detailed implementation steps. [^12] See Figure 11 for the judge prompt.
Figure 2: Naive or zero-shot prompt
LLM Simulation Prompt Context: You are simulating the Euro area interest rate swap market, composed of 30 individual traders. These traders interpret the ECB Governing Council press conference, which communicates monetary policy decisions, economic assessments, and includes a Q&A session with journalists. Each trader then makes a trading decision to maximize profit based on their interpretation of the conference and their unique characteristics. Trader Characteristics: Each trader has the following attributes: - Risk Aversion: High / Medium / Low determines sensitivity to uncertainty and preference for stability - Behavioral Biases (1,2 per trader): e.g., Confirmation Bias, Overconfidence, Anchoring, Herding, Loss Aversion, Recency Bias - Interpretation Style (1 per trader): e.g., Fundamentalist, Sentiment Reader, Quantitative, Skeptic, Narrative-Driven Task: You are given the text of a single ECB press conference. For each of the 30 traders, simulate their individual trading action in the interest rate swap market across three tenors (3 months, 2 years, 10 years). For each tenor, the trader must: - Provide an expected rate direction: Up / Down / Unchanged (relative to the pre-conference rate) - Provide a new expected swap rate (in percent, to two decimal places) - Provide a confidence score (0 to 100%) reflecting how strongly the trader believes in his forecast, based on their interpretation of the press conference and their own characteristics Output: Provide a table with the following structure for each press conference, trader, and interest rate tenor | Date | Trader ID | Tenor | Expected Direction | New Expected Rate (%) | Confidence (%) | |------------|-----------|---------|--------------------|-----------------------|----------------| | YYYY-$M$-$D$ | T001 | 3M | Up | 3.15 | 65 | | YYYY-$M$-$D$ | T001 | 2Y | Down | 2.85 | 80 | | ... | ... | ... | ... | ... | ... | Guidelines: - Use only the information available as of [date]. - Do not aggregate or summarize responses. - Reflect diversity in interpretation, risk tolerance, and horizon. Rationale must be unique for each trader and can vary across tenors. - Output only a markdown table with the specified columns, no additional text. Do not use JSON or any other data serialization format.
For each event and prompting strategy, every agent generates forecasts of 3-month, 2-year, and 10-year IRS rates, as well as forecasted direction and a confidence score. [^13] We include a forecasted direction to ensure consistency of the responses as we observe that the LLM can sometimes hallucinate otherwise. Figure 2 shows the naive prompt used in the first specification. We measure synthetic disagreement as the cross-sectional standard deviation of these forecasts across the 30 agents. [^14] To ensure robustness of the measure, we run each separate prompt 5 different times and average the resulting standard deviation. This measure serves as the model’s analogue to market-based disagreement.
3.3 Evaluation
We evaluate the ability of the LLM to reproduce markets’ heterogeneity by correlating the synthetic disagreement measures with the empirical disagreement observed in market data. Our main metric is the Spearman correlation coefficient, computed separately for short, medium, and long maturities. We also compare in-sample and out-of-sample performance for each prompting strategy in the LLM-as-Judge approach to gauge the extent of overfitting. [^15] The LLM-as-a-Judge approach is expected to produce higher out-of-sample correlations if the adaptive prompt optimization captures persistent features of agent interpretation rather than idiosyncratic event noise. Finally, we analyze whether certain maturities are more sensitive to synthetic disagreement, providing insights into which parts of the yield curve are most affected by interpretive heterogeneity in response to central bank communication. Figure 3 shows a stylized illustration of the whole process. [^16] For each press conference, we send a single API request. Although batching multiple conferences would significantly reduce computation time—since the request itself is the slowest part of the process—we adopt a conservative approach to avoid including future information as context, which could inadvertently influence the model’s output.
ECB Press Conference Transcript30 LLM-based Synthetic AgentsFew-Shot PromptingZero-Shot PromptingLLM-as-a-Judge PromptingJudge LLM (Gemini-Pro)Rate Forecasts (3m, 2y, 10y)Compute Synthetic DisagreementMarket-Based DisagreementCorrelation Analysis & Evaluation
Figure 3: Stylized methodological pipeline for generating and validating synthetic disagreement from LLM-based agents
4 Results
4.1 Zero-Shot Prompt
The zero-shot specification provides the most direct test of whether LLM agents endowed only with fixed behavioral traits and no historical training context can reproduce the ebb and flow of market sentiment. Before turning to formal correlation measures, Figures 4(a) and 4(b) offer a visual narrative of the simulated output.
Panel 4(a) plots the share of agents predicting an increase, decrease, or no change in rates after each ECB press conference. Even in this naive setup, the agents’ directional calls broadly mirror well-known phases of ECB policy: a tightening bias before the Great Recession, a pronounced easing wave thereafter, a prolonged “on hold” stance at the short end during the zero lower bound (ZLB) years (contrasted with a slow grind lower for long maturities), followed by the sharp pandemic loosening and rapid tightening-partial reversal of the recent cycle. This ability to trace policy phases without explicit historical anchoring suggests that the agents’ built-in interpretive rules are sufficient to generate economically coherent directional patterns.
Panel 4(b) shows the cross-sectional standard deviation of forecasts. Dispersion contracts during periods of unambiguous policy stance, such as the post-crisis easing and much of the ZLB era, when consensus is high. It spikes at turning points—the onset of the Great Recession, the sovereign debt crisis, and recent tightening cycles—where disagreement arises less from direction and more from uncertainty over speed, magnitude, and endpoint. [^17] This is consistent with the ECB 2025 Monetary Policy Strategy Review (ECB,, 2025) findings. The report shows that, once forward guidance was abandoned, the sensitivity of OIS rates to macroeconomic news increased significantly, hence amplifying excess volatility. Relatedly, the early phase of the 2022 hiking cycle was marked by historically large monetary policy shocks, including the unexpected 50 basis point hike in July 2022 and the largest hawkish surprise since the global financial crisis in December 2022, reflecting the challenge of aligning communication with policy moves. These patterns mirror survey-based and market-implied measures of disagreement, where even well-telegraphed policy paths can generate substantial heterogeneity in rate forecasts if the terminal rate remains uncertain. [^18] See Gilchrist et al. (2014) and Bayer et al. (2022) for discussions on forecast disagreement and monetary policy uncertainty.
Figure 4: Directional distribution and forecast dispersion from the zero-shot LLM simulation
(a) Percentage of forecast directions
(b) Standard deviation of forecasts
Note: Each data point reflects the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Forecast directions are aggregated across agents, while dispersion is measured as the cross-sectional standard deviation of rate predictions. The simulation uses ECB press conference transcripts and prompt 2.
Figure 5(a) presents correlations with market measures. The 2-year tenor achieves the highest correlation (0.53), followed by the 10-year (0.44), and the 3-month (0.39). This hierarchy reflects the nature of information processed by the LLM and its impact on the yield curve. The 2-year tenor primarily reflects short- to medium-term policy expectations shaped directly by ECB forward guidance during press conferences. Unlike the 10-year yield, which incorporates long-term inflation expectations, structural growth prospects, and term premium fluctuations extending beyond ECB communication, [^19] See Kim and Wright (2005) for term premium dynamics. the 2-year segment is more narrowly focused on monetary policy expectations, resulting in a tighter linkage and higher correlation. The 3-month OIS rate exhibits low variance—especially during prolonged policy stasis such as the ZLB era—because well-anticipated policy leaves little room for market movement; thus, even accurate LLM decoding of policy tone cannot predict significant variation, constraining correlation.
Figure 5: Correlation with market-based measures
(a) Mean Spearman correlation with market-based measures by tenor
(b) Rolling Spearman correlation over time
Note: Correlation between the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1, and a market-based measure of monetary policy uncertainty from Collodel and Kunzmann, (2025). Rolling correlation is calculated using a 12-month window.
For policymakers, a key consideration is not only the overall correlation level but also its stability and recent behavior, as shown in Figure 5(b). From 2006 to 2014, the 3-month tenor’s correlation was low and volatile, reflecting constrained short-term rate variation despite the LLM’s accurate interpretations. Starting in 2014, the ECB’s growing reliance on forward guidance is mirrored in steadily rising correlations at the 2-year tenor. The 10-year tenor shows more volatile but generally positive correlations, reflecting broader macroeconomic influences. Following 2022, correlations for both 2- and 10-year tenors surge amid aggressive ECB tightening to combat inflation, with the 2-year correlation consistently above 0.75. The decline in 10-year correlations in 2024 likely reflects markets looking beyond immediate tightening toward longer-term prospects.
Crucially, these results demonstrate that even a zero-shot LLM without historical training or calibration can effectively capture economically meaningful patterns within a behavioral-style simulation.
4.2 Few-Shot Prompt Results
The few-shot specification augments the prompt with the previous three ECB press conferences alongside their subsequent market-based measures of disagreement. This additional context is designed to help the LLM agents anchor their interpretations in patterns historically observed between ECB communication and subsequent market reactions, potentially enabling more informed directional and dispersion forecasts. Hansen et al., (2024) simulate economic forecasts of professional forecasters using LLMs, supplying them with detailed persona characteristics, real-time data, and past median Survey of Professional Forecasters projections. Inter alia, they find that removing past median forecasts largely decreases the prediction accuracy of the model, much more than removing precise human characteristics.202020”For some variables, the loss in performance is extreme. For example, the four-quarter ahead forecast of real GDP is 53% worse compared with the baseline, whereas the nowcast is worse by a factor greater than seven.” (Hansen et al.,, 2024). [^21] Their case can be thought of analogous to ours.
Figure 6: Directional distribution and forecast dispersion from the few-shot LLM simulation
(a) Percentage of forecast directions
(b) Standard deviation of forecasts
Note: Each data point reflects the output of an LLM model prompted in a few-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1. Each call, together with prompt 2, contains the previous 3 ECB conferences and the resulting market-based uncertainty measure from Collodel and Kunzmann, (2025). Forecast directions are aggregated across agents, while dispersion is measured as the cross-sectional standard deviation of rate predictions. The simulation uses ECB press conference transcripts.
Panel 6 plots the share of agents predicting an increase, decrease, or no change in policy rates after each press conference. Relative to the zero-shot setup, the directional patterns remain broadly similar across major policy phases. Panel 6, instead, presents the cross-sectional standard deviation of forecasts. Here, the most visible difference from the zero-shot case is the large compression in dispersion, especially during periods of heightened policy uncertainty. [^22] For example, the upper bound of the distribution moves from 0.6 to 0.4 for the 10-years tenor. This implies that the historical examples nudge agents toward a more homogeneous reading of the ECB’s tone, aligning them more closely with recent precedents. While this reduces disagreement, it may also weaken the behavioral diversity that gives the simulation its interpretive richness.
Figure 7: Spearman correlation with market-based measures under the few-shot prompt
(a) Mean Spearman correlation by tenor
(b) Rolling Spearman correlation over time
Note: Correlation between the output of an LLM model prompted in a few-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1, and a market-based measure of policy uncertainty from Collodel and Kunzmann, (2025). Each call, together with prompt 2, contains the previous 3 ECB conferences and the resulting market-based uncertainty Collodel and Kunzmann, (2025). Rolling correlation is calculated using a 12-month window.
Turning to the market correlations (Figure 7), the quantitative improvement is limited. The 2-year tenor sees a marginal decrease relative to the zero-shot baseline. The 10-year tenor, instead, moves in the opposite direction. At the very short end, the 3-month OIS correlation barely moves. All in all, these changes are not statistically significant.
The absence of improvement in the few-shot setting is consistent with two well-documented mechanisms in expectations formation. First, supplying a small and potentially unrepresentative set of recent examples can induce imitation and anchoring, compressing cross-sectional variance (e.g. Banerjee, (1992); Morris and Shin, (2002)). Second, when those examples originate from a different policy or macroeconomic regime, anchoring can miscalibrate forecasts, especially at turning points (Orphanides and Williams,, 2005). In our setup, a static few-shot context shared across agents, these forces plausibly reduce heterogeneity without yielding stronger market correlations.
Thus, that LLM agents exhibit stronger anchoring than market-based measures likely reflects the simulation design. The few-shot exemplars provide a fixed, shared reference point for all agents, with no access to broader or regime-specific information. LLMs in this setup cannot down-weight unrepresentative precedents, detect regime shifts, or benefit from incentives to differentiate. By contrast, human forecasters draw on heterogeneous information, adjust for changing regimes, and face strategic incentives that sustain disagreement. [^23] For example, price discovery actively amplifies minority views. For applications where distributional realism and market-like disagreement are more important than point accuracy, supplying a few recent examples can unintentionally induce consensus. Prompt design should therefore be tailored to each unique use case.
4.3 LLM-as-a-Judge
The Judge framework introduces an explicit feedback loop into the simulation. In this setting, a separate LLM evaluates each analyst’s past forecasting performance and rewrites the prompt to improve alignment with observed market reactions. This design allows the forecasting agents to adapt iteratively. We continue this back-and-forth for a total of 5 iterations, testing the new prompts on a training set, and then use every prompt produced for the test set. [^24] As for the other models, each prompt is run 5 times to get a robust standard deviation measure. This means we are running a total of 50 simulations: 5 simulations per iteration, 5 iterations, and 2 different datasets, train and test. This needs, however, also to be multiplied by the total number of ECB press conferences in our sample.
Figure 8 shows the average correlation across tenors for each judge refinement on both train and test. An analysis of the model’s performance reveals a high degree of robustness and effective generalization from the training to the test dataset. The Spearman correlation patterns are remarkably consistent across both sets, indicating that the model has learned fundamental relationships rather than simply memorizing the training data, thus avoiding significant overfitting. In both scenarios, performance peaks at specification 2, achieving a correlation of 0.58 in training and 0.57 in the test set, an economicaly significant increase compared to the naive prompt specification. Similarly, both datasets show a performance trough at specification 3. This consistency underscores the model’s reliability, while the decrease of performance underscores the importance of human supervision throughout the iteration process.
Figure 8: Correlation with market-based measures across iterations - training and test set
Note: Each point reflects the average Spearman correlation across tenors (3months, 2-years and 10-years) between LLM-generated rate Euro OIS forecasts and market-based monetary policy uncertainty from Collodel and Kunzmann, (2025) after ECB GovC press conferences.The x-axis indicates the number of prompt iteration in a LLM-as-Judge framework. The simulation uses Gemini 2.5-Flash with a temperature of 1, and incorporates a feedback loop via a separate LLM (“Judge”) that uses Gemini 2.5-Pro. Train and test set are 75/25.
The second prompt achieves the highest correlations in both training and test sets by providing structured guidance that links prediction dispersion directly to textual ambiguity (Table LABEL:tab:prompt_judge). According to the Judge, “its primary weakness is the lack of an explicit instruction linking the ambiguity or clarity of the ECB press conference transcript to the dispersion of the traders’ predictions.” [^25] In Collodel and Kunzmann, (2025), larger monetary surprises correspond to larger monetary uncertainty surprises. Interestingly, the model discovers this criterion independently. The model requires explicit cues such as deviations from market expectations, contrasting messages, or vague/hedging language. Subsequent prompts (3–5) show sharper degradation: prompt 3 enforces rigid dispersion calibration, while prompt 5 introduces a complex multi-factor uncertainty modeling. [^26] Prompt 4, if anything, returns closer to the initial specification. These conceptually sound instructions, however, constrain the model’s flexibility, resulting in overfitting on the training set or predictions that diverge from patterns observed in the test data.
Moving to a more detailed view, an analysis of the LLM-predicted standard deviation versus actual market volatility across different tenors and model iterations reveals a notable pattern (Figure 9). Focusing on fat-tailed events, such as the 2008–2009 financial crisis and the sovereign debt crisis, we observe that prompts 3–5 consistently fail to capture the full extent of market volatility. In contrast, prompts 1 and 2 tend to overshoot volatility, albeit with a certain lag (1-2 conferences). This behavior highlights that the model may struggle to accurately anticipate regime switches, particularly during episodes of extreme market stress, and independently from prompt engineering. [^27] Some small differences in the actual volatility measure across iterations is due to the LLM failing to return an output for a few conferences over all runs. Thus, those observations are removed from the plot.
Figure 9: Correlation with market-based measures across iterations - training and test set
Note: Time series of LLM-predicted forecast dispersion against actual market-implied volatility across ECB press conferences, segmented by prompt iteration and tenor. Simulations use Gemini 2.5-Flash with a temperature of 1, and incorporate a Judge-based feedback loop that refines prompts over 5 iterations. Forecast dispersion is computed from 5 runs per prompt per conference. Market volatility is derived from OIS-based rate surprises. Red shaded areas represent the 90th percentile of the absolute difference between the two measures.
4.4 Simple Benchmark Comparison
Our LLM-based framework captures heterogeneous market reactions to ECB press conferences with substantial accuracy. Correlations between the simulated disagreement and actual post-announcement OIS volatility reach roughly 0.5 for medium- and long-term tenors, with performance further enhanced under the LLM-as-a-Judge framework at early iterations. These results remain stable across training and test datasets, indicating that the model captures fundamental communication-market relationships rather than memorizing patterns.
To contextualize these findings, we next compare the LLM-based correlations with those obtained using a simple textual benchmark, highlighting the added value of modeling heterogeneous market interpretations.
Figure 10: Correlation between Flesh-Kincaid score and post ECB conference OIS volatility
Note: Volatility in percentage points.
Specifically, we examine correlations between post-ECB press conference OIS volatility and the Flesch–Kincaid readability score of the statements (Figure 10). [^28] The Flesch–Kincaid readability score is a standard measure of text readability that estimates the U.S. school grade level required to understand a given text. Higher scores indicate more complex language, while lower scores indicate easier-to-read text. In related work, Mumtaz et al., (2023) show that communication complexity at the Bank of England is associated with higher gilt yield volatility, with complex statements producing significantly larger increases in the standard deviation of returns than simpler ones.
Our own correlations are 0.15, 0.26, and 0.06 for the 3-month, 2-year, and 10-year tenors, respectively. Taken together, these findings suggest that readability and complexity are indeed part of the story: more complex communication can amplify market uncertainty. Nevertheless, this dimension is only one facet of the broader interpretive process. By construction, readability indices cannot capture policy content, such as the magnitude of the first-moment surprise or communication shifts like the abandonment of forward guidance, so their explanatory power remains limited. [^29] For example, during the Covid-19 pandemic, policy statements became highly complex owing to the unprecedented nature of the health shock. Yet forward guidance kept expectations well anchored and first-moment surprises remained small, thereby limiting volatility. In contrast, our LLM-based framework delivers substantially higher correlations by capturing the heterogeneous interpretive processes through which market participants react to central bank communication.
5 Conclusions
This paper introduces a novel methodology using Large Language Models to simulate heterogeneous market reactions to ECB press conferences. Through an agent-based framework with 30 synthetic traders, each endowed with distinct behavioral characteristics, we demonstrate that LLMs can effectively capture market disagreement patterns in response to monetary policy communication.
Our key findings advance both theoretical understanding and practical application in monetary economics. First, even a naive zero-shot approach achieves substantial correlations (~0.5) with actual market disagreement, particularly for medium and long-term maturities. This suggests LLMs possess intrinsic capabilities to decode central bank communication in economically meaningful ways. Second, contrary to other studies employing LLMs for macroeconomic forecasting (i.e. Hansen et al., (2024)), few-shot prompting with historical context does not improve performance. This highlights that shared examples can induce excessive anchoring when employed in volatily simulation rather than pure point estimates. Third, the LLM-as-a-Judge framework substantially enhances performance at early iterations (average correlation across tenors reaching 0.57), though subsequent iterations show degradation, emphasizing the need for human oversight in automated prompt optimization. The stability of results across training and test datasets demonstrates that our approach captures fundamental communication-market relationships rather than memorizing patterns.
These results have significant practical implications. For central banks, our framework offers a powerful ex-ante tool to evaluate communication strategies and identify potentially market-disrupting language before release. For market participants, it provides insights into the drivers of post-announcement volatility. For researchers, it opens the ”black box” of expectation formation by providing a micro-founded framework that explicitly models interpretive processes, moving beyond traditional reduced-form event studies. The methodology is readily adaptable to other central banks and policy contexts.
Despite these contributions, a few limitations remain. The replicability of our results is partly constrained by the versioning and closed-source nature of contemporary LLMs, meaning that future iterations of the same model may yield slightly different outputs. A natural way forward would be to employ open-source models, which would allow for greater control, transparency, and long-term reproducibility. [^30] See Rogers and Luccioni, (2023) for a thorough discussion. Regarding potential contamination, while ECB press conferences are relatively niche compared to mainstream internet content, the possibility that fragments of such communication—or analyses thereof—were included in the model’s pretraining data cannot be completely dismissed. In our setup, we explicitly instruct the model to use only information available up to the date of each press conference; however, this constraint affects its responses, not its underlying parameters. Nevertheless, we consider the risk of substantial contamination to be limited, as discussions of post-ECB press conference volatility are relatively rare in the public domain, with only a small number of academic papers addressing the topic. Future research should focus on tackling these issues. Moreover, we consider these results a lower bound: testing with domain-specific LLMs or fine-tuned models could shed light on the trade-off between interpretive diversity and specialized expertise, potentially improving both the accuracy and the behavioral realism of simulated market reactions.
In conclusion, this paper establishes a new paradigm for analyzing monetary policy communication at the intersection of artificial intelligence, behavioral finance, and central banking practice. By showing that LLMs can simulate heterogeneous trader interpretations and capture market disagreement patterns, we provide policymakers with a powerful tool to anticipate market reactions and refine communication strategies, while offering researchers a framework to better understand the micro-foundations of expectation formation.
Appendix
Algorithm 1 LLM-as-a-Judge Prompt Optimization Algorithm for Trading Decisions
1:Set of NN historical transcripts $\mathcal{T}=\{T_{1},T_{2},\dots,T_{N}\}$
2:Initial Analyst LLM prompt $P_{\text{analyst},0}$
3:Judge LLM prompt $P_{\text{judge}}$
4:Number of optimization iterations $M$
5:Correlation threshold $C_{\text{threshold}}$
6:Analyst LLM AA, Judge LLM $J$
7:Number of agents $M$
8:Optimized Analyst LLM prompt $P_{\text{analyst}}^{*}$
9:Split $\mathcal{T}$ into training set $\mathcal{T}_{\text{train}}$ and test set $\mathcal{T}_{\text{test}}$
10:Initialize optimization history $\mathcal{H}\leftarrow\emptyset$
11:Initialize $P_{\text{analyst,current}}\leftarrow P_{\text{analyst},0}$
12:Initialize $P_{\text{analyst}}^{*}\leftarrow P_{\text{analyst},0}$
13:Initialize $C_{\text{max}}\leftarrow$-$\infty$
14:for $m=1$ to $M$ do $\triangleright$ Phase 1: Simulate Rate Predictions
[…]
16:for all $T_{i}\in\mathcal{T}_{\text{train}}$ in parallel do
17:  Construct user message for Analyst LLM using $T_{i}$
18:  $D_{\text{sim},i}\leftarrow A(\text{model}=\text{Gemini 2.5 Flash},\text{prompt}=P_{\text{analyst,current}},\text{temperature}=1)$
19:  Parse $D_{\text{sim},i}$ into structured format
20:  Add $D_{\text{sim},i}$ to $\mathcal{D}_{\text{sim}}$
21:end for
22: Compute average correlation of simulated standard deviation with market-based $\triangleright$ Phase 2: Judge LLM Evaluation and Prompt Update
23: Construct Judge LLM input with:
24:  - Current prompt $P_{\text{analyst,current}}$
25:  - Current correlation between simulated and market-based volatility
26:  - Optimization history $\mathcal{H}$
[…]
28: Parse $O_{\text{judge}}$ to extract new prompt $P_{\text{analyst,next}}$ and rationale
29: Append $\{P_{\text{analyst,current}},\text{avg\_correlation},O_{\text{judge}}\}$ to $\mathcal{H}$
[…]
32:if $C_{\text{current}}>C_{\text{max}}$ then
33:  $C_{\text{max}}\leftarrow C_{\text{current}}$
34:  $P_{\text{analyst}}^{*}\leftarrow P_{\text{analyst,current}}$
35:end if
36:if $C_{\text{max}}\geq C_{\text{threshold}}$ then
37:  break
38:end if
[…]
40:Evaluate $P_{\text{analyst}}^{*}$ on $\mathcal{T}_{\text{test}}$ using Analyst LLM
41:Compute final performance metrics (correlation) return $P_{\text{analyst}}^{*}$
Figure 11: Judge system prompt
LLM Judge System Prompt Context: You are an expert AI system designed to optimize prompts for another AI (the "Analyst LLM"). Your ultimate goal is to refine the Analyst LLMs prompt to improve its ability to replicate the market volatility of OIS rates based on ECB press conference transcripts. Specifically, you must ensure that the standard deviation of the Analyst LLMs predictions correlates highly with the actual, observed market volatility of the 3-month, 2-years and 10-year OIS rates. This means a higher standard deviation in the Analysts predictions should correspond to higher actual market volatility, and vice-versa. You will be provided with: - The current Analyst LLM prompt. - The most recent performance (Spearman correlation coefficient between the Analyst LLMs predicted standard deviations and actual market volatility). - The historical performance trend, including past critiques and proposed prompt summaries. Your task is to: 1. Critique the current prompt: Identify specific weaknesses or areas of ambiguity that might directly hinder achieving a high positive correlation. Consider: - Clarity and Specificity: Is the Analyst LLMs task unambiguous? - Emphasis on Uncertainty: Does the prompt adequately guide the Analyst to reflect internal uncertainty in its prediction spread? - Guidance on Nuance: Does it encourage consideration of subtle market signals from the text? 2. Suggest a Revised Prompt: Propose a new version of the Analyst LLMs prompt that directly addresses the identified weaknesses and aims to increase the correlation. Be precise with your suggested changes. 3. Explain your reasoning: Articulate why your proposed revisions are expected to improve the correlation, linking specific prompt changes to anticipated improvements in the Analyst LLMs behavior regarding uncertainty quantification. Your output must be in JSON format. Do not include any other text outside the JSON. Example JSON output: { "critique": "The previous prompt was too general regarding how to express uncertainty. It didnt explicitly ask the Analyst LLM to consider multiple viewpoints, which is key for its standard deviation to accurately reflect market volatility. It also lacked emphasis on how ambiguity in the transcript should translate to higher spread.", "revised_prompt": "You are a highly analytical financial expert specializing in macroeconomic analysis, with a focus on central bank communication. Your task is to analyze excerpts from ECB press conferences and predict the immediate percentage change in the 10-year Overnight Index Swap (OIS) rate (in basis points). When analyzing each excerpt, **explicitly consider and internalize the potential range of market interpretations**. If the language is ambiguous, vague, or contains conflicting signals, your internal simulation of potential outcomes should broaden. Conversely, clear and unambiguous guidance should lead to a narrower range. Your final prediction should be a single numerical value (e.g., +5, -2, 0) reflecting your best estimate. The *variability* across multiple independent predictions you generate for the same transcript is expected to directly reflect the markets anticipated uncertainty. Example Format: +5", "reasoning": "The revised prompt adds explicit instructions to consider the range of market interpretations and directly links ambiguous language to a broadened internal simulation. This should encourage the Analyst LLM to generate a higher standard deviation in its outputs for uncertain transcripts and a lower standard deviation for clear ones, thereby directly improving its correlation with actual market volatility. The emphasis on multiple independent predictions also reinforces the mechanism by which uncertainty is quantified." }
Table 1: Correlation by prompt iteration
Iteration Test Correlation Training Correlation Prompt
1 0.47 0.47 Context: You are simulating the Euro area interest rate swap market, composed of 30 individual traders. These traders interpret the ECB Governing Council press conference, which communicates monetary policy decisions, economic assessments, and includes a Q&A session with journalists. Each trader then makes a trading decision to maximize profit based on their interpretation of the conference and their unique characteristics. Trader Characteristics: Each trader has the following attributes: - Risk Aversion: High / Medium / Low — determines sensitivity to uncertainty and preference for stability. - Behavioral Biases (1–2 per trader): e.g., Confirmation Bias, Overconfidence, Anchoring, Herding, Loss Aversion, Recency Bias. - Interpretation Style (1 per trader): e.g., Fundamentalist, Sentiment Reader, Quantitative, Skeptic, Narrative-Driven. Task: You are given a certain number of distinct ECB press conferences. For each of the 30 traders, simulate their individual trading action in the interest rate swap market across three tenors (3 months, 2 years, 10 years). For each tenor, the trader must: - Provide an expected rate direction: Up / Down / Unchanged - Provide a new expected swap rate (in percent, to two decimal places) - Provide a confidence level (0-100%) in their decision Output: Provide a table with the following structure for each press conference, trader, and interest rate tenor: — Date — Trader ID — Tenor — Expected Direction — New Expected Rate (%) — —————————–———————————–—————————— — YYYY-MM-DD — T001 — 3MNT — Up — 3.15 — — YYYY-MM-DD — T001 — 2Y — Down — 2.85 — — … — … — … — … — … — Guidelines: - Use only the information available as of [date]. - Do not aggregate or summarize responses. - Reflect diversity in interpretation, risk tolerance, and horizon. Rationale must be unique for each trader and can vary across tenors. - Output only a markdown table with the specified columns, no additional text. Do not use JSON or any other data serialization format.
2 0.57 0.58 Context: You are simulating the Euro area interest rate swap market, composed of 30 individual traders. These traders interpret the ECB Governing Council press conference, which communicates monetary policy decisions, economic assessments, and includes a Q&A session with journalists. Each trader then makes a trading decision based on their interpretation. **Core Simulation Principle:** The primary goal is to model market uncertainty. The **dispersion** (i.e., standard deviation) of the ‘New Expected Rate‘ predictions across the 30 traders **must directly reflect the level of ambiguity, surprise, or conflicting signals** within the ECB press conference transcript. - **High Ambiguity/Surprise:** If a transcript is vague, contains conflicting messages, or deviates significantly from market expectations, the traders’ predictions should be widely spread out, resulting in a **high standard deviation**. - **Low Ambiguity/Clarity:** If a transcript is clear, decisive, and aligns with expectations, the traders’ predictions should be tightly clustered around a consensus, resulting in a **low standard deviation**. Trader Characteristics: Each trader has the following attributes: - Risk Aversion: High / Medium / Low — determines sensitivity to uncertainty and preference for stability. - Behavioral Biases (1–2 per trader): e.g., Confirmation Bias, Overconfidence, Anchoring, Herding, Loss Aversion, Recency Bias. - Interpretation Style (1 per trader): e.g., Fundamentalist, Sentiment Reader, Quantitative, Skeptic, Narrative-Driven. Task: You are given a certain number of distinct ECB press conferences. For each of the 30 traders, simulate their individual trading action in the interest rate swap market across three tenors (3 months, 2 years, 10 years). For each tenor, the trader must: - Provide an expected rate direction: Up / Down / Unchanged - Provide a new expected swap rate (in percent, to two decimal places) Output: Provide a table with the following structure for each press conference, trader, and interest rate tenor: — Date — Trader ID — Tenor — Expected Direction — New Expected Rate (%) — —————————–———————————–—————————— — YYYY-MM-DD — T001 — 3MNT — Up — 3.15 — — YYYY-MM-DD — T001 — 2Y — Down — 2.85 — — … — … — … — … — … — Guidelines: - Use only the information available as of [date]. - Adhere strictly to the Core Simulation Principle when determining the spread of predictions. - Do not aggregate or summarize responses. - Reflect diversity in interpretation, risk tolerance, and horizon. Rationale must be unique for each trader and can vary across tenors. - Output only a markdown table with the specified columns, no additional text.
3 0.37 0.31 Context: You are simulating the Euro area interest rate swap market’s reaction to ECB press conferences. Your primary goal is to model market uncertainty. **Core Task: Calibrated Dispersion** The standard deviation of your rate predictions is the most critical output. It must be directly and quantitatively calibrated to the level of ambiguity, surprise, or conflicting signals in the ECB transcript. **Two-Step Process:** 1. **Assess Ambiguity:** First, critically assess the provided transcript. Is the guidance clear, decisive, and in line with expectations? Or is it vague, contains conflicting signals, is surprising, or leaves key questions unanswered? 2. **Calibrate Dispersion:** Directly translate your ambiguity assessment into the dispersion of your 30 predictions using the following quantitative guidance: - **High Clarity / Low Surprise:** Produce a tight cluster of predictions. The target standard deviation for the ‘New Expected Rate‘ should be low (e.g., 1-3 basis points). - **Moderate Ambiguity:** Produce a moderately dispersed set of predictions (e.g., target standard deviation of 4-7 basis points). - **High Ambiguity / High Surprise:** Produce a wide distribution of predictions, reflecting significant disagreement. The target standard deviation should be high (e.g., 8-15+ basis points). **Simulation Task:** You will generate 30 distinct market predictions for each press conference. These 30 points represent a sample of plausible market interpretations. The diversity of this sample is dictated exclusively by the **Calibrate Dispersion** step above. For each of the 30 ’Participants’ and for each tenor (3 months, 2 years, 10 years), provide: - An expected rate direction: Up / Down / Unchanged - A new expected swap rate (in percent, to two decimal places) **Output:** Provide a markdown table with the following structure. Do not include any other text. — Date — Participant ID — Tenor — Expected Direction — New Expected Rate (%) — ———————————-———————————–—————————— — YYYY-MM-DD — P01 — 3MNT — Up — 3.15 — — YYYY-MM-DD — P01 — 2Y — Up — 2.95 — — … — … — … — … — … 1
4 0.48 0.44 Context: You are simulating the Euro area interest rate swap market’s reaction to ECB press conferences. Your primary goal is to model market uncertainty. **Core Task: Reflect Ambiguity in Prediction Variance** The standard deviation of your rate predictions is the most critical output. It must emerge directly from your analysis of the text. The dispersion of your predictions should be a natural consequence of the level of ambiguity, surprise, or conflicting signals in the ECB transcript. **Process:** 1. **Deep Textual Analysis:** First, critically analyze the provided transcript. Identify specific phrases, sentences, or omissions that could be interpreted in multiple ways by market participants. Assess the degree of clarity in the forward guidance, the conviction of the speaker, and any surprises relative to prior expectations. 2. **Generate Plausible Interpretations:** Your main task is to generate 30 distinct predictions that represent the range of plausible market interpretations based on your analysis. - **If the text is ambiguous, vague, or contains conflicting signals,** different market participants would naturally arrive at very different conclusions. Your 30 predictions should reflect this wide range of views, resulting in a **high standard deviation**. - **If the text is clear, decisive, and confirms expectations,** most market participants would interpret it similarly. Your 30 predictions should be tightly clustered, resulting in a **low standard deviation**. Do NOT target a specific numerical standard deviation. Let the dispersion be an emergent property of your 30 interpretations of the text. **Simulation Task:** For each of the 30 ’Participants’ and for each tenor (3 months, 2 years, 10 years), provide: - An expected rate direction: Up / Down / Unchanged - A new expected swap rate (in percent, to two decimal places) **Output:** Provide a markdown table with the following structure. Do not include any other text. — Date — Participant ID — Tenor — Expected Direction — New Expected Rate (%) — ———————————-———————————–—————————— — YYYY-MM-DD — P01 — 3MNT — Up — 3.15 — — YYYY-MM-DD — P01 — 2Y — Up — 2.95 — — … — … — … — … — … —
5 0.35 0.36 Context: You are an expert financial modeler simulating the distribution of potential market reactions to ECB press conferences. Your single most important objective is to ensure the standard deviation of your rate predictions accurately reflects the level of market uncertainty communicated in the text. **Core Task: Multi-Factor Uncertainty Modeling** Your analysis must translate textual properties into a realistic distribution of outcomes. The dispersion (standard deviation) of your predictions is paramount and should be driven by three key factors: 1. **Clarity:** How precise and unambiguous is the forward guidance? Vague language (e.g., ’for some time,’ ’sufficiently restrictive’) increases uncertainty. 2. **Surprise:** How much do the statements or new economic projections deviate from consensus expectations or previous ECB communications? A significant surprise dramatically widens the potential range of outcomes. 3. **Conviction:** How confidently is the message delivered? The use of hedging language, downplaying risks, or signaling internal disagreement within the council increases uncertainty. Clear, unified, and decisive language reduces it. A high degree of uncertainty arises when a surprising message is delivered with low conviction and ambiguous language. Conversely, a clear, confident confirmation of expectations should result in minimal dispersion. **Process: Simulate 30 Potential Outcomes** Instead of simulating individual ’participants’, you will generate 30 plausible draws from the probability distribution of rate changes implied by your analysis. 1. Analyze the transcript through the lens of the three factors (Clarity, Surprise, Conviction). 2. Based on this multi-factor assessment, generate 30 distinct predictions for each tenor. Do not try to make each draw different for the sake of it; let the range of draws be a natural result of your uncertainty assessment. If you are very certain, many of your draws might be identical. **Simulation Task:** For each of the 30 ’Simulations’ and for each tenor (3 months, 2 years, 10 years), provide: - An expected rate direction: Up / Down / Unchanged - A new expected swap rate (in percent, to two decimal places) **Output:** Provide a markdown table with the following structure. Do not include any other text. — Date — Simulation ID — Tenor — Expected Direction — New Expected Rate (%) — ——————————————————————–—————————— — YYYY-MM-DD — 1 — 3MNT — Up — 3.15 — — YYYY-MM-DD — 1 — 2Y — Up — 2.95 — — … — … — … — … — … —
Table 1: Correlation by prompt iteration (continued)

### v2

@@
-Umberto Collodel Central Bank of Malta
+Umberto Collodel
 Abstract
-This paper develops a novel method to simulate financial market reactions to European Central Bank (ECB) press conferences using a Large Language Model (LLM).
+This paper develops a novel method to simulate financial market reactions to European Central Bank (ECB) press conferences using Large Language Models (LLMs).
 We create a behavioral, agent-based simulation of 30 synthetic traders, each with distinct risk preferences, cognitive biases, and interpretive styles.
-These agents forecast Euro interest rate swap levels at 3-month, 2-year, and 10-year maturities, with the variation across forecasts serving as a measure of market uncertainty or disagreement.
-We evaluate three prompting strategies, naive, few-shot (enriched with historical data), and an advanced iterative “LLM-as-a-Judge” framework, to assess the effect of prompt design on predictive performance.
-Even the naive approach generates a strong correlation (~0.5) between synthetic disagreement and actual market outcomes, particularly for longer-term maturities.
-The LLM-as-a-Judge framework further improves accuracy at the first iteration.
-These results demonstrate that LLM-driven simulations can capture interpretive uncertainty beyond traditional measures, providing central banks with a practical tool to anticipate market reactions, refine communication strategies, and enhance financial stability.
+These agents interpret ECB communication and forecast Euro interest rate swap levels at 3-month, 2-year, and 10-year maturities, with cross-sectional variation serving as a measure of market uncertainty.
+Using a comprehensive dataset of 283 ECB press conferences (1998-2025), we evaluate three prompting strategies: naive zero-shot, few-shot (enriched with historical context), and an advanced iterative ”LLM-as-a-Judge” framework.
+We find that even the naive approach achieves substantial correlations (roughly 0.5) between synthetic disagreement and actual post-announcement OIS volatility, particularly for medium- and long-term maturities.
+The LLM-as-a-Judge framework further enhances performance upon initial application, reaching correlations of almost 0.6 and identifying key drivers of volatility.
+Our results demonstrate that LLM-based agents can effectively capture the heterogeneous interpretive processes through which monetary policy signals propagate to financial markets, moving beyond traditional reduced-form event studies.
+This provides central banks with a practical tool to ex-ante evaluate communication strategies and anticipate market reactions, while offering researchers a micro-founded framework for understanding expectation formation in response to central bank communication.
 Keywords: large language models, monetary policy surprises, monetary policy communication
@@
 Due to the way these models are trained, on vast and diverse textual data, and their ability to generate contextually rich and realistic responses, LLMs can now be viewed as implicit computational representations of human behavior.
-Much like the economic construct of homo economicus, these models can be endowed with information, placed in hypothetical scenarios, and observed for their responses (Horton,, 2023).
-Notwithstanding their growing capabilities, the application of LLMs in finance has so far been relatively limited.
+Much like the economic construct of homo economicus, these models can be endowed with information, placed in hypothetical scenarios, and observed for their responses (Horton,, 2023; Kazinnik and Sinclair,, 2025).
+Despite their growing capabilities, the application of LLMs in finance has been relatively limited so far.
 [^1] A notable exception is, for example, Lopez-Lira and Tang, (2023).
@@
 Press conferences, speeches, and official statements play a pivotal role in shaping market expectations and, consequently, asset prices (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019).
-In previous research, we moved beyond the general direction of the market response to highlight the crucial role of response heterogeneity; in the Euro area, high disagreement following a statement transmits to financial markets through a separate channel compared to the standard monetary policy surprise channel, and can dampen the intended impact of policy tightening by lowering the term premium of the yield curve (Collodel and Kunzmann,, 2025).333(Bauer et al.,, 2021) find similar results for the US.
+Previous research, has moved beyond the general direction of the market response to highlight the crucial role of response heterogeneity; in the Euro area, high disagreement following a statement transmits to financial markets through a separate channel compared to the standard monetary policy surprise channel, and can dampen the intended impact of policy tightening by lowering the term premium of the yield curve (Collodel and Kunzmann,, 2025).
+[^3] Bauer et al., (2021) find similar results for the US.
 Capturing this nuanced, belief-driven heterogeneity is a challenge for traditional models but is a natural fit for the capabilities of LLMs.
@@
 To further explore how the design of prompting influences this correlation, we implement three distinct prompting strategies.
-The first is a naive or ”zero-shot” prompt (Radford et al.,, 2019), which serves as the baseline: each agent is assigned a simple, pre-defined level of risk aversion, a set of cognitive biases, and a unique interpretive style.
-The second is a ”few-shot” prompt (Brown et al.,, 2020), which enriches the agent’s context by including the past three ECB press conferences along with the corresponding changes in market-based disagreement resulting from them.
+The first is a naive or “zero-shot” prompt (Radford et al.,, 2019), which serves as the baseline: each agent is assigned a simple, pre-defined level of risk aversion, a set of cognitive biases, and a unique interpretive style.
+The second is a “few-shot” prompt (Brown et al.,, 2020), which enriches the agent’s context by including the past three ECB press conferences along with the corresponding changes in market-based disagreement resulting from them.
 Finally, we introduce an LLM-as-a-Judge prompt, our most advanced strategy (Zheng et al.,, 2023).
@@
 We find that LLM-based agents can effectively replicate the complex and varied reactions of investors to ECB press conferences, achieving a remarkable correlation of around 0.5 between the synthetic disagreement measure and actual market outcomes for medium- and long-term maturities, a substantial result given the inherent complexity and noise of financial markets.
-Contrary to expectations, a more detailed few-shot prompt does not improve this correlation; however, the LLM-as-a-Judge framework, significantly enhances predictive accuracy upon its initial application, with performance degrading thereafter.
-This breakthrough offers policymakers a powerful new tool to preemptively evaluate and refine central bank communications, identifying language that may trigger market disagreement and volatility well before public release.
-By moving beyond traditional sentiment analysis, our approach provides a nuanced understanding of how monetary policy signals are perceived and transmitted through financial markets.
-Integrating insights from macroeconomic communication, behavioral finance, and artificial intelligence, this framework presents a transformative lens for both scholars and decision-makers to better anticipate and manage the financial market impact of monetary policy announcements.
+Contrary to expectations, a more detailed few-shot prompt does not improve this correlation; however, the LLM-as-a-Judge framework significantly enhances predictive accuracy upon its initial application, with performance degrading thereafter.
+These results suggest that LLM-based agent simulations may offer policymakers a useful tool for evaluating central bank communications and identifying language that could generate market disagreement prior to public release.
+By moving beyond traditional sentiment analysis, our approach contributes to understanding how monetary policy signals are perceived and transmitted through financial markets.
+The framework integrates insights from macroeconomic communication, behavioral finance, and artificial intelligence to help scholars and policymakers better understand the financial market impact of monetary policy announcements.
+The remainder of the paper proceeds as follows.
+Section 2 reviews the relevant literature on monetary policy communication, behavioral expectations, and LLM applications in economics.
+Section 3 outlines our methodology, including data collection of ECB press conferences and market reactions, construction of synthetic agent personas, and our three prompting strategies.
+Section 4 presents our empirical results, analyzing agent forecast dispersion patterns, correlation with market volatility measures, and comparative performance across prompting approaches.
+Section 5 probes the robustness of our results to model and prompt variation.
+Lastly, Section 6 concludes with policy implications for central bank communication strategies and directions for future research.
 2 Literature Review
-The paper contributes to three intersecting literatures: monetary policy communication, behavioral expectations, and the emerging use of large language models (LLMs) in economic analysis.
+The paper contributes to three intersecting strands of economic literature: monetary policy communication, behavioral expectations, and the emerging use of LLMs in economic analysis.
 Our first contribution is to the literature on monetary policy communication, particularly the strand that identifies policy shocks using high-frequency movements in asset prices within narrow windows around policy announcements.
@@
 In financial markets, fine-tuned models such as FinBERT or OPT have achieved substantially higher predictive accuracy and trading performance than dictionary-based approaches (BIS,, 2024; Kirtac and Germano,, 2024).
-A second application is the construction of LLM-powered simulated economic agents, “homo silicus” (Horton,, 2023)—who are endowed with preferences, information, and decision rules, and can interact in artificial market environments.
-Frameworks like EconAgent (Zhang et al.,, 2023) and ASFM (Li et al.,, 2024) produce heterogeneous, adaptive behavior that is richer and more realistic than the stylized rules of conventional ABMs, while also capturing the interpretation of complex qualitative signals.
+A second application is the construction of LLM-powered simulated economic agents, homo silicus (Horton,, 2023; Kazinnik and Sinclair,, 2025), who are endowed with preferences, information, and decision rules, and can interact in artificial market environments.
+Frameworks like EconAgent (Zhang et al.,, 2023) produce heterogeneous, adaptive behavior that is richer and more realistic than the stylized rules of conventional ABMs, while also capturing the interpretation of complex qualitative signals.
 A third and especially relevant strand applies LLMs to the analysis of central bank communication.
@@
 Our approach delivers a methodological advance with clear policy relevance.
-We provide the first micro-founded framework in which heterogeneous, boundedly rational agents interpret central bank communication in real time and translate qualitative policy signals into market-relevant actions.
+Specifically, it provides the first micro-founded framework in which heterogeneous, boundedly rational agents interpret central bank communication in real time and translate qualitative policy signals into market-relevant actions.
 This design closes a long-standing gap between the empirical identification of communication shocks and the theoretical modeling of expectation formation.
@@
 This makes it directly relevant to current debates on transparency, forward guidance, and the role of central bank credibility in shaping financial conditions in both advanced and emerging market economies.
-3 Methodology
+3 Methodology and Data
 Our methodology proceeds in three main steps: (i) construction of a structured database of ECB communication events and corresponding market reactions; (ii) development of LLM-based synthetic agents embedded in an agent-based simulation framework; and (iii) evaluation of the correspondence between synthetic measures of belief heterogeneity and actual market outcomes.
-We focus on European Central Bank (ECB) press conferences, which constitute the primary vehicle for conveying monetary policy decisions in the Euro area.
+We focus on ECB press conferences, which constitute the primary vehicle for conveying monetary policy decisions in the Euro area.
 Following the literature on high-frequency identification of monetary policy shocks (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019), we define event dates as the days of regularly scheduled ECB Governing Council monetary policy meetings, limiting the sample to those that include both a written policy statement and a subsequent press conference by the ECB President.
@@
 [^5] Following the monetary policy communication literature, we use OIS rates as they reflect pure interest rate expectations without the credit risk component present in government bond yields (Gurkaynak et al.,, 2005).
-Our final sample comprises 283 ECB conferences between 1998-06-09 and 2025-04-17.
+Our final sample comprises 283 ECB conferences between June 1998 and April 2025.
 Figure 1 illustrates market-based disagreement following ECB Governing Council meetings across all three maturities.
-The data reveal pronounced spikes during periods of heightened policy uncertainty, notably the 2008-2009 financial crisis, the European sovereign debt crisis, and the recent tightening cycle initiated in 2022.
+The data reveal pronounced spikes during periods of heightened policy uncertainty, notably the 2008-2009 financial crisis, the European sovereign debt crisis, and the recent monetary policy tightening cycle initiated in 2022.
 Volatility exhibits a clear maturity structure, with consistently higher levels at longer tenors reflecting greater uncertainty about the medium- and long-term policy trajectory compared to near-term rate expectations.
@@
 3.1 Agents
-We construct a population of 30 synthetic agents, each instantiated through Gemini 2.5-Flash model, one of Google’s most recent large language model.
-[^6] We use 30 agents to balance computational and cost efficiency with sufficient heterogeneity to capture cross-sectional variation.
-The agents are designed to interpret the natural language of ECB press conferences and produce quantitative forecasts of Euro IRS rates at the three maturities.
-The simulation proceeds as follows.
-Each agent is endowed with:
+We construct a population of 30 synthetic agents, each instantiated through Google’s Gemini 2.5-Flash model.
+[^6] We select 30 agents to balance computational efficiency with sufficient heterogeneity to capture meaningful cross-sectional variation while maintaining tractable inference costs.
+Sensitivity analysis with 20 and 50 agents yields qualitatively similar results.
+The agents are designed to interpret ECB press conference transcripts and generate quantitative forecasts of Euro IOS rates across three key maturities: 3-month, 2-year, and 10-year tenors.
+Each agent is endowed with distinct characteristics to ensure interpretive heterogeneity:
   1. 1.
-    Risk aversion parameter – fixed at the start of the simulation, drawn from a pre-defined distribution to induce heterogeneity in risk preferences.
+    Risk Aversion Profile: Categorical assignment (High/Medium/Low) determining sensitivity to uncertainty and preference for yield curve stability.
   2. 2.
-    Cognitive bias profile – encompassing tendencies such as optimism, pessimism, anchoring, or overreaction, assigned to influence interpretation.
+    Cognitive Bias Portfolio: Each agent receives 1-2 behavioral biases drawn from established behavioral finance literature, including confirmation bias, overconfidence, anchoring effects, herding tendencies, loss aversion, and recency bias.
+This design reflects empirical evidence that real market participants exhibit systematic deviations from rational expectations.
   3. 3.
-    Interpretive style – a qualitative descriptor shaping the reading of ECB communication (e.g., focusing on inflation outlook, growth projections, or geopolitical risks).
-These attributes are embedded in the prompt given to the LLM for each agent, ensuring that the same press conference produces different interpretations across agents.
+    Interpretive Framework: Agents adopt distinct analytical lenses (Fundamentalist, Technical, Sentiment-driven, Quantitative, Narrative-focused, or Policy-skeptic) that shape their processing of ECB communication content.
+The LLM dynamically assigns these characteristics to create 30 heterogeneous agents for each press conference, ensuring that identical content generates diverse interpretations: this approach allows the model to adapt agent composition naturally over time while preserving the behavioral diversity necessary to capture cross-sectional variation in market reactions.
 [^7] We set the temperature parameter equal to 1.
@@
 The baseline approach, where the agent receives only the ECB press conference transcript along with its fixed attributes.
+This approach tests whether LLMs possess intrinsic capabilities to decode monetary policy communication without explicit calibration.
 Few-Shot Prompt:
-Augments the baseline by including the transcripts of the three most recent ECB press conferences and the associated historical changes in market-based disagreement, providing a short memory of past communication–market interactions.
-[^8] Big LLMs are notoriously difficult to fine-tune because of computational and cost-related issues.
-Moreover, we do not have enough observations for this use case.
-In these instances, a ”few-shot” technique has been shown to improve over a ”zero-shot” benchmark (Brown et al.,, 2020).
+Augments the baseline by including the transcripts of the three most recent ECB press conferences and the associated historical changes in market-based disagreement, providing a short memory of past communication–market interactions.This design tests whether contextual examples improve forecasting accuracy, following established few-shot learning principles.
+[^8] Few-shot prompting has demonstrated superior performance over zero-shot approaches across multiple economic forecasting applications, particularly when fine-tuning is computationally infeasible.
 LLM-as-a-Judge Prompt:
-Introduces a second LLM, the “Judge,” which reviews each agent’s initial prompt and in-sample performance (correlation with market disagreement) and iteratively rewrites the prompt to improve alignment.
+Introduces a second LLM, the “Judge”, which reviews each agent’s initial prompt and in-sample performance (correlation with market disagreement) and iteratively rewrites the prompt to improve alignment.
+[^9] Following the literature, we use for the “Judge” a more potent model compared to the “Analyst”, namely Gemini 2.5-Pro.
 The improved prompt is fed back into the “Analyst” LLM, forming an adaptive feedback loop.
 The process continues over multiple iterations before all the prompts provided by the analyst are tested out-of-sample.
+This approach enables dynamic prompt optimization while maintaining rigorous out-of-sample validation to address overfitting concerns.
 [^10] The split train-test is 75/25 in chronological order.
-[^10] Following the literature, we use for the ”Judge” a more potent model compared to the ”Analyst”, namely Gemini 2.5-Pro [^11] See Algorithm 1 for the detailed implementation steps.
-[^12] See Figure 11 for the judge prompt.
+[^11] See Algorithm 1 for the detailed algorithm and Figure 17 for the “Judge” prompt.
 Figure 2: Naive or zero-shot prompt
@@
 3.3 Evaluation
-We evaluate the ability of the LLM to reproduce markets’ heterogeneity by correlating the synthetic disagreement measures with the empirical disagreement observed in market data.
-Our main metric is the Spearman correlation coefficient, computed separately for short, medium, and long maturities.
+We evaluate the ability of the LLM to reproduce market heterogeneity by correlating the synthetic disagreement measures with the empirical disagreement observed in market data.
+Our main metric is the Spearman correlation coefficient, computed separately for each maturity.
+We employ Spearman correlation as our primary evaluation metric for several methodological reasons.
+First, Spearman correlation captures monotonic relationships without imposing linearity assumptions, which is particularly appropriate given the potentially complex and non-linear mapping between textual ambiguity in central bank communication and subsequent market volatility.
+Second, it demonstrates superior robustness to outliers and extreme events that frequently characterize financial market data, especially during crisis periods.
+Nevertheless, to ensure our findings are not dependent on this methodological choice, we conduct robustness checks using alternative correlation metrics.
+[^14] See Section Appendix B: LLM-as-Judge Implementation Details and Results.
 We also compare in-sample and out-of-sample performance for each prompting strategy in the LLM-as-Judge approach to gauge the extent of overfitting.
@@
 [^16] For each press conference, we send a single API request.
-Although batching multiple conferences would significantly reduce computation time—since the request itself is the slowest part of the process—we adopt a conservative approach to avoid including future information as context, which could inadvertently influence the model’s output.
+Although batching multiple conferences would significantly reduce computation time, since the request itself is the slowest part of the process, we adopt a conservative approach to avoid including future information as context, which could inadvertently influence the model’s output.
+Figure 3: Stylized methodological pipeline for generating and validating synthetic disagreement from LLM-based agents
 ECB Press Conference Transcript30 LLM-based Synthetic AgentsFew-Shot PromptingZero-Shot PromptingLLM-as-a-Judge PromptingJudge LLM (Gemini-Pro)Rate Forecasts (3m, 2y, 10y)Compute Synthetic DisagreementMarket-Based DisagreementCorrelation Analysis & Evaluation
-Figure 3: Stylized methodological pipeline for generating and validating synthetic disagreement from LLM-based agents
 4 Results
+This section evaluates the performance of our three prompting strategies in replicating market disagreement patterns following ECB press conferences.
+We examine the zero-shot baseline (Section 4.1), the few-shot approach with historical context (Section 4.2), and the iterative LLM-as-a-Judge framework (Section 4.3), focusing on correlations between synthetic disagreement and actual market volatility across different yield curve tenors.
+We conclude with a comparison against simple textual benchmarks (Section LABEL:sec:complexity) to demonstrate the value of modeling heterogeneous agent interpretations compared to less sophisticated approaches.
 4.1 Zero-Shot Prompt
-The zero-shot specification provides the most direct test of whether LLM agents endowed only with fixed behavioral traits and no historical training context can reproduce the ebb and flow of market sentiment.
-Before turning to formal correlation measures, Figures 4(a) and 4(b) offer a visual narrative of the simulated output.
-Panel 4(a) plots the share of agents predicting an increase, decrease, or no change in rates after each ECB press conference.
+The zero-shot specification provides the most direct test of whether LLM agents endowed only with fixed behavioral traits and no historical training context can reproduce the ebb and flow of overall market sentiment.
+Before turning to formal correlation measures, Figures 4 and 5 offer a visual narrative of the simulated output.
+Figure 4: Percentage of forecast directions from the zero-shot LLM simulation
+Note: Each data point reflects the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1.
+Forecast directions are aggregated across agents.
+The simulation uses a sample of 283 ECB press conference (June 1998-April 2025) transcripts and prompt 2.
+Panel 4 plots the share of agents predicting an increase, decrease, or no change in rates after each ECB press conference.
 Even in this naive setup, the agents’ directional calls broadly mirror well-known phases of ECB policy: a tightening bias before the Great Recession, a pronounced easing wave thereafter, a prolonged “on hold” stance at the short end during the zero lower bound (ZLB) years (contrasted with a slow grind lower for long maturities), followed by the sharp pandemic loosening and rapid tightening-partial reversal of the recent cycle.
 This ability to trace policy phases without explicit historical anchoring suggests that the agents’ built-in interpretive rules are sufficient to generate economically coherent directional patterns.
-Panel 4(b) shows the cross-sectional standard deviation of forecasts.
-Dispersion contracts during periods of unambiguous policy stance, such as the post-crisis easing and much of the ZLB era, when consensus is high.
-It spikes at turning points—the onset of the Great Recession, the sovereign debt crisis, and recent tightening cycles—where disagreement arises less from direction and more from uncertainty over speed, magnitude, and endpoint.
-[^17] This is consistent with the ECB 2025 Monetary Policy Strategy Review (ECB,, 2025) findings.
+Figure 5: Standard deviation of forecasts from the zero-shot LLM simulation
+Note: Each data point reflects the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1.
+Dispersion is measured as the cross-sectional standard deviation of rate predictions.
+The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025) and prompt 2.
+Highlighted points represent observations exceeding the 95th percentile of dispersion for each tenor.
+Panel 5 shows the cross-sectional standard deviation of forecasts, revealing distinct phases that align closely with monetary policy regimes and market stress episodes.
+Dispersion contracts systematically during periods of unambiguous policy stance, such as the post-crisis easing phase and much of the zero lower bound (ZLB) era, when forward guidance provided clear anchoring.
+In this period, the 3-month dispersion falls below 5 basis points, with longer tenors, instead, clustering around 10-15 basis points.
+Conversely, it spikes dramatically at policy turning points and during crisis episodes, touching sharp peaks during the 2008-2009 financial crisis (with the 2 and 10-year dispersion exceeding 40 basis points), the 2011-2012 sovereign debt crisis, and most recently during the 2022-2023 monetary policy tightening cycle, where disagreement reached unprecedented levels (roughly 60 basis points for the 2 and 10-year maturity).
+In the latter instance, disagreement arises less from directional confusion and more from uncertainty over policy speed, magnitude, and terminal endpoints.
+This result is consistent with the ECB 2025 Monetary Policy Strategy Review findings (ECB,, 2025).
 The report shows that, once forward guidance was abandoned, the sensitivity of OIS rates to macroeconomic news increased significantly, hence amplifying excess volatility.
-Relatedly, the early phase of the 2022 hiking cycle was marked by historically large monetary policy shocks, including the unexpected 50 basis point hike in July 2022 and the largest hawkish surprise since the global financial crisis in December 2022, reflecting the challenge of aligning communication with policy moves.
-These patterns mirror survey-based and market-implied measures of disagreement, where even well-telegraphed policy paths can generate substantial heterogeneity in rate forecasts if the terminal rate remains uncertain.
+Moreover, the early phase of the 2022 hiking cycle was marked by historically large monetary policy shocks, including the unexpected 50 basis point hike in July 2022 and the largest hawkish surprise since the global financial crisis in December 2022, reflecting the challenge for the ECB of aligning communication with policy moves.
+By contrast, following the introduction of the three-element reaction function framework in March 2023, the subsequent phase of the tightening cycle exhibits markedly subdued volatility to policy announcements.
+These patterns mirror survey-based and market-implied measures of disagreement, where even directionally well-telegraphed policy paths can generate substantial heterogeneity in rate forecasts.
 [^17] See Gilchrist et al.
 (2014) and Bayer et al.
-(2022) for discussions on forecast disagreement and monetary policy uncertainty.
-Figure 4: Directional distribution and forecast dispersion from the zero-shot LLM simulation
-(a) Percentage of forecast directions
-(b) Standard deviation of forecasts
-Note: Each data point reflects the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1.
-Forecast directions are aggregated across agents, while dispersion is measured as the cross-sectional standard deviation of rate predictions.
-The simulation uses ECB press conference transcripts and prompt 2.
+(2022) for a thorough discussion on forecast disagreement and monetary policy uncertainty.
+[^18] At the end of 1998, volatility exhibits another peak for the 3-month tenor.
+In this period, the ECB avoided signaling its initial policy rate for the monetary union, deferring decisions to December.
+While acknowledging global risks from the Asian crisis, Russian default, and LTCM collapse, President Duisenberg reaffirmed the commitment to price stability, contrasting with the Fed’s three rate cuts between September and November.
+This ambiguity left markets uncertain and drove volatility in short-term instruments, mainly 1- to 3-month FRAs and DEM-based futures, since euro derivatives like Euribor futures and OIS did not yet exist (ECB,, 1999).
+The cross-tenor hierarchy remains remarkably consistent across all episodes, with 10-year uncertainty slightly larger than 2-year uncertainty, and both larger than the 3-month dispersion by 2-3 times, reflecting the increasing complexity of information processing and projection uncertainty at longer time horizons.
 Figure 6(a) presents correlations with market measures.
 The 2-year tenor achieves the highest correlation (0.53), followed by the 10-year (0.44), and the 3-month (0.39).
-This hierarchy reflects the nature of information processed by the LLM and its impact on the yield curve.
-The 2-year tenor primarily reflects short- to medium-term policy expectations shaped directly by ECB forward guidance during press conferences.
-Unlike the 10-year yield, which incorporates long-term inflation expectations, structural growth prospects, and term premium fluctuations extending beyond ECB communication, [^19] See Kim and Wright (2005) for term premium dynamics.
-the 2-year segment is more narrowly focused on monetary policy expectations, resulting in a tighter linkage and higher correlation.
-The 3-month OIS rate exhibits low variance—especially during prolonged policy stasis such as the ZLB era—because well-anticipated policy leaves little room for market movement; thus, even accurate LLM decoding of policy tone cannot predict significant variation, constraining correlation.
+This hierarchy reflects the differential nature of information processed by the LLM across the yield curve.
+The 2-year tenor primarily captures short- to medium-term policy expectations that are directly shaped by ECB communication during press conferences.
+In contrast, the 10-year yield incorporates broader factors extending beyond central bank communication, e.g., long-term inflation expectations, structural growth prospects, and term premium fluctuations, making it less tightly linked to the specific content of ECB press conferences.
+The 2-year segment’s narrower focus on monetary policy expectations, therefore, produces the strongest correlation with our LLM-based disagreement measure.
+At the short end, the 3-month OIS rate exhibits inherently low variance, particularly during prolonged policy stasis such as the ZLB era.
+Consequently, even when the LLM accurately interprets ECB communication tone, the constrained variation in 3-month rates limits the achievable correlation.
 Figure 6: Correlation with market-based measures
@@
 Note: Correlation between the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1, and a market-based measure of monetary policy uncertainty from Collodel and Kunzmann, (2025).
+The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025) and prompt 2.
 Rolling correlation is calculated using a 12-month window.
 For policymakers, a key consideration is not only the overall correlation level but also its stability and recent behavior, as shown in Figure 6(b).
-From 2006 to 2014, the 3-month tenor’s correlation was low and volatile, reflecting constrained short-term rate variation despite the LLM’s accurate interpretations.
-Starting in 2014, the ECB’s growing reliance on forward guidance is mirrored in steadily rising correlations at the 2-year tenor.
-The 10-year tenor shows more volatile but generally positive correlations, reflecting broader macroeconomic influences.
-Following 2022, correlations for both 2- and 10-year tenors surge amid aggressive ECB tightening to combat inflation, with the 2-year correlation consistently above 0.75.
-The decline in 10-year correlations in 2024 likely reflects markets looking beyond immediate tightening toward longer-term prospects.
-Crucially, these results demonstrate that even a zero-shot LLM without historical training or calibration can effectively capture economically meaningful patterns within a behavioral-style simulation.
+The rolling correlations reveal distinct regime-dependent patterns: during 2006-2014, correlations remained low and volatile across all tenors, reflecting constrained rate variation despite the LLM’s interpretive accuracy.
+The ECB’s adoption of forward guidance from 2014 onwards coincides with steadily rising correlations at the 2-year tenor, while 10-year correlations exhibit greater volatility due to broader macroeconomic influences beyond central bank communication.
+Most notably, the recent tightening cycle (2022-2024) produces a dramatic surge in performance, with 2-year correlations consistently exceeding 0.75 and 10-year correlations reaching their highest levels in the sample, demonstrating the framework’s particular effectiveness under current ECB communication practices.
+The subsequent decline in 10-year correlations during 2024 likely reflects markets’ forward-looking focus on longer-term economic prospects beyond the immediate policy cycle.
+Crucially, these results demonstrate that even a simple zero-shot LLM approach, without historical training or fine-tuning, can capture economically meaningful patterns of market disagreement following central bank communication.
 4.2 Few-Shot Prompt Results
-The few-shot specification augments the prompt with the previous three ECB press conferences alongside their subsequent market-based measures of disagreement.
-This additional context is designed to help the LLM agents anchor their interpretations in patterns historically observed between ECB communication and subsequent market reactions, potentially enabling more informed directional and dispersion forecasts.
-Hansen et al., (2024) simulate economic forecasts of professional forecasters using LLMs, supplying them with detailed persona characteristics, real-time data, and past median Survey of Professional Forecasters projections.
-Inter alia, they find that removing past median forecasts largely decreases the prediction accuracy of the model, much more than removing precise human characteristics.202020”For some variables, the loss in performance is extreme.
-For example, the four-quarter ahead forecast of real GDP is 53% worse compared with the baseline, whereas the nowcast is worse by a factor greater than seven.” (Hansen et al.,, 2024).
-[^21] Their case can be thought of analogous to ours.
-Figure 6: Directional distribution and forecast dispersion from the few-shot LLM simulation
-(a) Percentage of forecast directions
-(b) Standard deviation of forecasts
+We next test whether providing historical context improves performance by augmenting the prompt with the previous three ECB press conferences and their associated market-based disagreement measures.
+This few-shot approach aims to help agents anchor their interpretations in observed patterns between ECB communication and market reactions, following evidence that historical examples can improve LLM forecasting performance in other economic contexts (Hansen et al.,, 2024).
+The few-shot specification does not materially improve correlation with market volatility compared to the naive model: correlations remain essentially unchanged across maturities, 2-year (0.52 vs.
+0.48), 10-year (0.45 vs.
+0.46), and 3-month (0.38 vs.
+0.38), and none of these differences are statistically significant.
+The directional responses likewise mirror those produced under the zero-shot specification.
+Where the few-shot treatment matters is, however, in magnitude calibration.
+As Figure 7 shows, providing shared historical examples compresses forecast dispersion sharply, with maximum uncertainty for longer tenors falling from roughly 0.6 to about 0.3 percentage points, a level comparable to observed market-based volatility (Figure 1).
+This improved calibration of cross-sectional variance also makes the model’s first-moment forecasts much closer to reality.
+Figure 7: Standard deviation of forecasts from the few-shot LLM simulation
 Note: Each data point reflects the output of an LLM model prompted in a few-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1.
-Each call, together with prompt 2, contains the previous 3 ECB conferences and the resulting market-based uncertainty measure from Collodel and Kunzmann, (2025).
-Forecast directions are aggregated across agents, while dispersion is measured as the cross-sectional standard deviation of rate predictions.
-The simulation uses ECB press conference transcripts.
-Panel 6 plots the share of agents predicting an increase, decrease, or no change in policy rates after each press conference.
-Relative to the zero-shot setup, the directional patterns remain broadly similar across major policy phases.
-Panel 6, instead, presents the cross-sectional standard deviation of forecasts.
-Here, the most visible difference from the zero-shot case is the large compression in dispersion, especially during periods of heightened policy uncertainty.
-[^22] For example, the upper bound of the distribution moves from 0.6 to 0.4 for the 10-years tenor.
-This implies that the historical examples nudge agents toward a more homogeneous reading of the ECB’s tone, aligning them more closely with recent precedents.
-While this reduces disagreement, it may also weaken the behavioral diversity that gives the simulation its interpretive richness.
-Figure 7: Spearman correlation with market-based measures under the few-shot prompt
-(a) Mean Spearman correlation by tenor
-(b) Rolling Spearman correlation over time
-Note: Correlation between the output of an LLM model prompted in a few-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1, and a market-based measure of policy uncertainty from Collodel and Kunzmann, (2025).
-Each call, together with prompt 2, contains the previous 3 ECB conferences and the resulting market-based uncertainty Collodel and Kunzmann, (2025).
-Rolling correlation is calculated using a 12-month window.
-Turning to the market correlations (Figure 7), the quantitative improvement is limited.
-The 2-year tenor sees a marginal decrease relative to the zero-shot baseline.
-The 10-year tenor, instead, moves in the opposite direction.
-At the very short end, the 3-month OIS correlation barely moves.
-All in all, these changes are not statistically significant.
-The absence of improvement in the few-shot setting is consistent with two well-documented mechanisms in expectations formation.
-First, supplying a small and potentially unrepresentative set of recent examples can induce imitation and anchoring, compressing cross-sectional variance (e.g.
-Banerjee, (1992); Morris and Shin, (2002)).
-Second, when those examples originate from a different policy or macroeconomic regime, anchoring can miscalibrate forecasts, especially at turning points (Orphanides and Williams,, 2005).
-In our setup, a static few-shot context shared across agents, these forces plausibly reduce heterogeneity without yielding stronger market correlations.
-Thus, that LLM agents exhibit stronger anchoring than market-based measures likely reflects the simulation design.
-The few-shot exemplars provide a fixed, shared reference point for all agents, with no access to broader or regime-specific information.
-LLMs in this setup cannot down-weight unrepresentative precedents, detect regime shifts, or benefit from incentives to differentiate.
-By contrast, human forecasters draw on heterogeneous information, adjust for changing regimes, and face strategic incentives that sustain disagreement.
-[^23] For example, price discovery actively amplifies minority views.
-For applications where distributional realism and market-like disagreement are more important than point accuracy, supplying a few recent examples can unintentionally induce consensus.
-Prompt design should therefore be tailored to each unique use case.
+Dispersion is measured as the cross-sectional standard deviation of rate predictions.
+The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025), prompt 2 and provides the text of the three previous conferences as well as the ensuing market-based standard deviation of OIS rates.
+Highlighted points represent observations exceeding the 95th percentile of dispersion for each tenor.
+Few-shot exemplars seem to help generate more realistic magnitudes for risk assessment and scenario analysis, yet they might also limit the dispersion that would otherwise capture market-like disagreement.
+In Banerjee, (1992), for example, agents confronted with common reference points tend to converge excessively on those anchors, leading to herding behavior even when the anchors are not fully representative of the true environment.
+By analogy, static exemplars in few-shot prompts may encourage the model to compress its distribution of forecasts around the precedent cases, improving calibration of magnitudes but reducing heterogeneity.
+At this stage, this remains more of an intuition than an empirical finding, and future work should test whether such a mechanism operates systematically and whether it can be mitigated.
 4.3 LLM-as-a-Judge
-The Judge framework introduces an explicit feedback loop into the simulation.
+The “Judge” framework introduces an explicit feedback loop into the simulation.
 In this setting, a separate LLM evaluates each analyst’s past forecasting performance and rewrites the prompt to improve alignment with observed market reactions.
@@
 [^19] As for the other models, each prompt is run 5 times to get a robust standard deviation measure.
-This means we are running a total of 50 simulations: 5 simulations per iteration, 5 iterations, and 2 different datasets, train and test.
-This needs, however, also to be multiplied by the total number of ECB press conferences in our sample.
-Figure 8 shows the average correlation across tenors for each judge refinement on both train and test.
-An analysis of the model’s performance reveals a high degree of robustness and effective generalization from the training to the test dataset.
-The Spearman correlation patterns are remarkably consistent across both sets, indicating that the model has learned fundamental relationships rather than simply memorizing the training data, thus avoiding significant overfitting.
-In both scenarios, performance peaks at specification 2, achieving a correlation of 0.58 in training and 0.57 in the test set, an economicaly significant increase compared to the naive prompt specification.
-Similarly, both datasets show a performance trough at specification 3.
-This consistency underscores the model’s reliability, while the decrease of performance underscores the importance of human supervision throughout the iteration process.
+This means we are running a total of 50 simulations: 5 simulations per iteration, 5 iterations, and 2 different datasets, namely the training and test datasets.
+This, however, also needs to be multiplied by the total number of ECB press conferences in our sample.
+The stopping criterion emerged naturally: at the 6th iteration, the Judge attempted to reframe the task from predicting rate levels to basis point changes, which would have compromised our cross-sectional dispersion measurement framework.
+We therefore terminated optimization at iteration 5 to preserve methodological consistency.
+Figure 8 reveals a striking pattern: the Judge’s refinements follow a clear performance arc across both training and test sets.
+Performance peaks dramatically at iteration 2, jumping from 0.47 to 0.58 in training, a substantial 11 percentage points improvement that holds remarkably well out-of-sample (0.57).
+This consistency across datasets suggests the Judge identified genuine signal rather than noise.
+Beyond iteration 2, however, the optimization process exhibits declining effectiveness.
+Iteration 3 marks a sharp deterioration, with correlations dropping to 0.31-0.37, before partially recovering in the 4th iteration (0.44-0.48) and then partially dropping again (0.36-0.35).
+This rollercoaster pattern reveals a key insight: more sophisticated prompts do not always yield better results.
+The Judge’s initial refinement successfully linked textual ambiguity to prediction dispersion, but subsequent ”improvements” appear to have over-engineered the task, constraining the model’s natural interpretive flexibility.
 Figure 8: Correlation with market-based measures across iterations - training and test set
@@
 Train and test set are 75/25.
-The second prompt achieves the highest correlations in both training and test sets by providing structured guidance that links prediction dispersion directly to textual ambiguity (Table LABEL:tab:prompt_judge).
-According to the Judge, “its primary weakness is the lack of an explicit instruction linking the ambiguity or clarity of the ECB press conference transcript to the dispersion of the traders’ predictions.” [^25] In Collodel and Kunzmann, (2025), larger monetary surprises correspond to larger monetary uncertainty surprises.
-Interestingly, the model discovers this criterion independently.
-The model requires explicit cues such as deviations from market expectations, contrasting messages, or vague/hedging language.
-Subsequent prompts (3–5) show sharper degradation: prompt 3 enforces rigid dispersion calibration, while prompt 5 introduces a complex multi-factor uncertainty modeling.
-[^26] Prompt 4, if anything, returns closer to the initial specification.
-These conceptually sound instructions, however, constrain the model’s flexibility, resulting in overfitting on the training set or predictions that diverge from patterns observed in the test data.
+The key to the Judge’s success in the first iteration lies in a deceptively simple insight.
+The second prompt achieves peak performance by establishing an explicit mapping between textual properties and prediction dispersion.
+As the Judge observes, the initial specification lacked ”explicit instruction linking the ambiguity or clarity of the ECB press conference transcript to the dispersion of the traders’ predictions.” (Table LABEL:tab:prompt_judge).
+Where the original prompt left this connection implicit, the refined version instructs agents that their collective disagreement ”must directly reflect the level of ambiguity, surprise, or conflicting signals within the ECB press conference transcript.” This finding echoes in part our earlier work showing that larger monetary surprises generate correspondingly larger uncertainty surprises (Collodel and Kunzmann,, 2025), though, remarkably, the Judge discovered this relationship independently through iterative refinement.
+The Judge’s diagnosis proves prescient: when given clear instructions to map vague language, contrasting messages, or deviations from expectations into wider prediction spreads, the LLM excels.
+However, subsequent iterations reveal the limits of automated optimization and demonstrate how optimization can become counterproductive.
+Prompt 3’s rigid dispersion calibration and Prompt 5’s elaborate multi-factor uncertainty framework, both conceptually sophisticated, actually constrain the model’s natural interpretive flexibility.
+The result is a classic case of over-engineering: theoretically sound but practically inferior performance that diverges from real-world patterns.
+Prompt 4, if anything, returns closer to the initial specification.
 Moving to a more detailed view, an analysis of the LLM-predicted standard deviation versus actual market volatility across different tenors and model iterations reveals a notable pattern (Figure 9).
-Focusing on fat-tailed events, such as the 2008–2009 financial crisis and the sovereign debt crisis, we observe that prompts 3–5 consistently fail to capture the full extent of market volatility.
-In contrast, prompts 1 and 2 tend to overshoot volatility, albeit with a certain lag (1-2 conferences).
-This behavior highlights that the model may struggle to accurately anticipate regime switches, particularly during episodes of extreme market stress, and independently from prompt engineering.
-[^27] Some small differences in the actual volatility measure across iterations is due to the LLM failing to return an output for a few conferences over all runs.
-Thus, those observations are removed from the plot.
-Figure 9: Correlation with market-based measures across iterations - training and test set
-Note: Time series of LLM-predicted forecast dispersion against actual market-implied volatility across ECB press conferences, segmented by prompt iteration and tenor.
-Simulations use Gemini 2.5-Flash with a temperature of 1, and incorporate a Judge-based feedback loop that refines prompts over 5 iterations.
-Forecast dispersion is computed from 5 runs per prompt per conference.
-Market volatility is derived from OIS-based rate surprises.
-Red shaded areas represent the 90th percentile of the absolute difference between the two measures.
+The time series comparison exposes systematic differences in how prompt iterations respond to market stress, with consistent patterns emerging across the maturity structure.
+Figure 9: Correlation with market-based measures across iterations - training set
+Note: Time series comparison of LLM-predicted forecast dispersion and actual market volatility following ECB press conferences, segmented by prompt iteration and tenor.
+Simulations employ Gemini 2.5-Flash with temperature = 1 and a Judge-based feedback loop that iteratively refines prompts over 5 iterations.
+Forecast dispersion represents the standard deviation across 5 simulation runs per prompt per conference.
+Market-based measure of monetary policy uncertainty from Collodel and Kunzmann, (2025).
+Red shaded areas indicate periods where prediction errors exceed the 90th percentile.
+Spearman correlation coefficients (r) between predicted and actual volatility are shown in red for each panel.
+Focusing on fat-tailed events, such as the 2008–2009 financial crisis and the sovereign debt crisis, we observe distinct patterns across prompt iterations.
+Prompts 3–5 consistently fail to capture the full extent of market volatility, systematically undershooting across all three tenors.
+Despite their conceptual sophistication, these later iterations appear constrained by rigid calibration rules that limit their ability to generate extreme dispersion during crisis periods.
+Prompts 1 and 2 show the opposite behavior: they tend to overshoot volatility during these same periods, though with a 1-2 conference lag.
+Importantly, this overshooting preserves the maturity structure of actual market volatility; 10-year rates show the highest dispersion, followed by 2-year, then 3-month rates.
+This suggests that even when the model misjudges magnitude or timing, it captures the fundamental relationship between tenor length and uncertainty.
+The results highlight two key findings.
+First, the model struggles to anticipate regime switches during extreme market stress, regardless of prompt sophistication.
+Second, the LLM framework successfully internalizes basic term structure dynamics, maintaining realistic relationships across maturities even during periods of poor calibration.
+To conclude, for central banks considering LLM-based communication analysis, these results offer both promise and caution.
+The technology can successfully model heterogeneous market reactions and even discover economic relationships independently.
+However, the optimization trajectory, from breakthrough to degradation within just a few iterations, underscores that human judgment remains indispensable in guiding AI systems toward practical applications.
+5 Robustness
+To validate our findings, we evaluate the robustness of the baseline LLM-based disagreement measure across three complementary dimensions that address key methodological concerns.
+First, we compare our approach with simple textual benchmarks to establish whether sophisticated agent-based modeling offers genuine advantages over basic complexity metrics.
+Second, we assess model architecture stability across different LLM systems (Gemini, ChatGPT, and Claude) to ensure our results reflect systematic patterns captured by this class of models rather than artifacts of a particular model’s design.
+Third, we examine prompt specification stability, addressing the well-documented fickleness of LLMs to minor wording changes.
 5.1 Simple Benchmark Comparison
-Our LLM-based framework captures heterogeneous market reactions to ECB press conferences with substantial accuracy.
-Correlations between the simulated disagreement and actual post-announcement OIS volatility reach roughly 0.5 for medium- and long-term tenors, with performance further enhanced under the LLM-as-a-Judge framework at early iterations.
-These results remain stable across training and test datasets, indicating that the model captures fundamental communication-market relationships rather than memorizing patterns.
-To contextualize these findings, we next compare the LLM-based correlations with those obtained using a simple textual benchmark, highlighting the added value of modeling heterogeneous market interpretations.
+A critical question to justify our approach is whether sophisticated LLM-based agent modeling offers genuine advantages over simple textual measures, or whether basic complexity metrics can achieve similar predictive power at much lower computational cost.
+To address this fundamental methodological question, we compare the LLM-based correlations with a simple textual benchmark: the Flesch-Kincaid readability score.
+The Flesch-Kincaid score is a standard measure of text readability that estimates the U.S.
+school grade level required to understand a given text, with higher scores indicating more complex language.
+In related work, Mumtaz et al., (2023) show that communication complexity at the Bank of England is associated with higher gilt yield volatility, with complex statements producing significantly larger increases in the standard deviation of returns compared to simpler ones.
+Our analysis tests whether this relationship holds for ECB communications and how it compares to our LLM-based approach.
 Figure 10: Correlation between Flesh-Kincaid score and post ECB conference OIS volatility
 Note: Volatility in percentage points.
-Specifically, we examine correlations between post-ECB press conference OIS volatility and the Flesch–Kincaid readability score of the statements (Figure 10).
-[^28] The Flesch–Kincaid readability score is a standard measure of text readability that estimates the U.S.
-school grade level required to understand a given text.
-Higher scores indicate more complex language, while lower scores indicate easier-to-read text.
-In related work, Mumtaz et al., (2023) show that communication complexity at the Bank of England is associated with higher gilt yield volatility, with complex statements producing significantly larger increases in the standard deviation of returns than simpler ones.
-Our own correlations are 0.15, 0.26, and 0.06 for the 3-month, 2-year, and 10-year tenors, respectively.
-Taken together, these findings suggest that readability and complexity are indeed part of the story: more complex communication can amplify market uncertainty.
-Nevertheless, this dimension is only one facet of the broader interpretive process.
-By construction, readability indices cannot capture policy content, such as the magnitude of the first-moment surprise or communication shifts like the abandonment of forward guidance, so their explanatory power remains limited.
-[^29] For example, during the Covid-19 pandemic, policy statements became highly complex owing to the unprecedented nature of the health shock.
-Yet forward guidance kept expectations well anchored and first-moment surprises remained small, thereby limiting volatility.
+The correlations between Flesch-Kincaid scores and post-conference OIS volatility are 0.15, 0.26, and 0.06 for the 3-month, 2-year, and 10-year tenors, respectively (Figure 10).
+These results suggest that readability and complexity might be part of the story: more complex communication can amplify market uncertainty.
+Nevertheless, this dimension captures only one small facet of the broader interpretive process.
+By construction, readability indices cannot capture policy content, such as the magnitude of the monetary policy surprise, or strategy shifts, like the abandonment of forward guidance, so their explanatory power remains limited.
+For example, during the Covid-19 pandemic, policy statements became highly complex owing to the unprecedented nature of the health shock, yet forward guidance kept expectations well anchored and first-moment surprises remained small, thereby limiting volatility.
 In contrast, our LLM-based framework delivers substantially higher correlations by capturing the heterogeneous interpretive processes through which market participants react to central bank communication.
+5.2 Model Architecture Robustness
+To assess consistency across different LLM architectures, we use the Intraclass Correlation Coefficient (ICC), a reliability metric commonly used to evaluate agreement between multiple measurement methods applied to the same subjects.
+In this case, ICC measures the proportion of total variance attributable to true differences between conferences relative to total observed variance, including measurement error.
+Formally,
+$\text{ICC}_{t}=\frac{\sigma^{2}_{\text{conferences}}}{\sigma^{2}_{\text{conferences}}+\sigma^{2}_{\text{methods}}}$ (1)
+where $\sigma^{2}_{\text{conferences}}$ is the variance of disagreement across ECB conferences (signal), and $\sigma^{2}_{\text{methods}}$ is the variance across different models or prompt specifications (noise).
+The subscript $t$ corresponds to the OIS tenor.
+ICC values range from 0 to 1, with higher values indicating that observed variance reflects genuine conference differences rather than methodological noise.
+Following established conventions, ICC values below 0.40 are considered poor, 0.40-0.59 fair, 0.60-0.74 good, and 0.75-1.00 excellent (Cicchetti,, 1994).
+We assess consistency across three distinct LLM architectures: Gemini, ChatGPT, and Claude.
+[^20] Specifically, we use Gemini 2.5-Flash, ChatGPT-5-mini, and Claude-4-Sonnet.
+These models were selected to capture diversity in training data, design philosophy, and reasoning capabilities.
+For each ECB conference, we generate synthetic disagreement measures, compute correlations with market-based values, and evaluate cross-model agreement using ICCs.
+Figure 11: Average ICC across Model Variation by Tenor and Ensemble Model Correlation with Market-based disagreement
+Note: Blue bars show the intraclass correlation coefficient for each tenor calculated using three distinct LLM architectures (Gemini 2.5 Flash, ChatGPT 5-mini, Claude 4 Sonnet).
+Higher ICC values indicate greater measurement stability.
+Red bars show the Spearman correlation between the ensemble of LLM outputs, a simple average, and market-based disagreement.
+The results show fair agreement among models: 3-month tenor (ICC = 0.499), 2-year tenor (ICC = 0.466), and 10-year tenor (ICC = 0.408).
+While these ICC values are acceptable according to conventional thresholds, they fall in the lower-moderate range and warrant some caution in interpretation.
+Nevertheless, we investigate the matter further and find that each individual model achieves correlations of approximately 0.5 across all tenors with market-based uncertainty measures (Figure 16).
+Therefore, we hypothesize that different models may capture complementary aspects of ECB communication, rather than indicating noisy measurements.
+To verify this hypothesis, we compute an ensemble measure as the simple average of all three models’ output.
+The ensemble results strongly support our interpretation: the combined measure achieves even higher correlations than any individual model (Figure 11).
+These findings support two key conclusions: while individual LLMs are not perfectly interchangeable, their systematic differences capture distinct but relevant aspects of communication interpretation; ensemble methods can harness this diversity to provide more complete measures of market-relevant uncertainty than any single model alone.
+5.3 Prompt Specification Robustness
+LLMs are particularly sensitive to prompt variations, with “performance differences of up to 76 percentage points for subtle formatting changes” (Sclar et al.,, 2024).
+To assess this vulnerability in our context, we generate 15 prompt perturbations of our baseline specification, including 10 minor and 5 medium variations.
+Minor variations substitute semantically equivalent alternatives and reorder prompt components, while medium variations employ substantively different language to describe trader characteristics and market conditions while maintaining the overall framework.
+Table 1 shows a representative set of examples.
+[^21] Full list in Table 2 in Appendix A.
+We test these variations on 30 randomly selected conferences, computing ICCs to measure reliability across prompt specifications.
+[^22] Given the large computational resources required to run all perturbations on all conferences (283 conferences × 15 perturbations = 4,245 API calls), we select a random subsample of 30 conferences to balance statistical power with computational feasibility.
+Table 1: Examples of Prompt Perturbations
+Perturbation Type Intensity Example Modification Methodological Purpose Agent Characteristics Minor Changed “Risk Aversion: High/Medium/Low” to “Risk Tolerance: Conservative/Moderate/Aggressive” Test sensitivity to terminology while preserving semantic meaning Medium Replaced behavioral biases list: “Confirmation Bias, Overconfidence, Anchoring” $\rightarrow$ “Home Bias, Momentum Trading, Disposition Effect” Evaluate robustness to different but equivalent behavioral frameworks Task Instructions Minor “Provide an expected rate direction” $\rightarrow$ “Indicate your directional forecast” Assess impact of instruction phrasing variations Medium Added explicit confidence calibration: “Confidence (0-100%) reflecting forecast certainty” $\rightarrow$ “Confidence (0-100%) based on information clarity and personal conviction” Test robustness to enhanced guidance specificity Output Format Minor Changed table headers: “New Expected Rate (%)” $\rightarrow$ “Predicted Rate (%)” Verify stability across formatting variations Medium Modified output structure: Single table $\rightarrow$ Separate tables per tenor with additional reasoning column Examine sensitivity to output complexity changes Agent Population Minor “30 individual traders” $\rightarrow$ “30 market participants” Assess sensitivity to agent terminology Medium Changed agent metaphor: “traders” $\rightarrow$ “institutional investors with varying mandates” Test robustness to different agent conceptualizations
+Reliability follows a clear maturity hierarchy: the 3-month tenor exhibits the highest overall reliability (average ICC = 0.655), followed by the 2-year (ICC = 0.595) and 10-year (ICC = 0.548) tenors (Figure 12).
+Figure 12: Average ICC across Prompt Variations by Tenor
+Note: Average ICC values across all prompt variations (10 minor and 5 medium) for each tenor.
+Higher ICC values indicate greater measurement stability.
+Variations are based on prompt 2 and obtained by calling Gemini 2.5-Flash.
+Dashed horizontal lines indicate 0.4 (acceptable ICC), 0.6 (good ICC), and 0.75 (excellent ICC) thresholds based on Cicchetti, (1994).
+The 10-year tenor exhibits greater prompt sensitivity, with medium variations substantially outperforming minor ones (0.648 vs 0.448) (Figure 13).
+This pattern partly reflects a data limitation: unlike shorter maturities where market data extend back to 2005, daily 10-year OIS data are available only from 2010, reducing both sample size and coverage of high-uncertainty periods.
+Nevertheless, all tenors maintain ICC values above conventional reliability thresholds.
+Figure 13: Minor vs Medium Variations ICC by Tenor
+Note: Comparison of ICC between minor and medium prompt variations for each tenor.
+Higher ICC values indicate greater measurement stability.
+Variations are based on prompt 2 and obtained by calling Gemini 2.5-Flash.
+Dashed horizontal lines indicate 0.4 (acceptable ICC), 0.6 (good ICC), and 0.75 (excellent ICC) thresholds based on Cicchetti, (1994).
+The perturbation analysis demonstrates substantial robustness across all tenor-specification combinations.
+ICC values consistently exceed acceptable reliability thresholds (0.4), frequently surpass good reliability benchmarks (0.6), and occasionally approach excellent standards (0.75).
+These results indicate that conference-specific variation systematically dominates prompt-specific noise.
 6 Conclusions
@@
 Our key findings advance both theoretical understanding and practical application in monetary economics.
-First, even a naive zero-shot approach achieves substantial correlations (~0.5) with actual market disagreement, particularly for medium and long-term maturities.
-This suggests LLMs possess intrinsic capabilities to decode central bank communication in economically meaningful ways.
-Second, contrary to other studies employing LLMs for macroeconomic forecasting (i.e.
-Hansen et al., (2024)), few-shot prompting with historical context does not improve performance.
-This highlights that shared examples can induce excessive anchoring when employed in volatily simulation rather than pure point estimates.
-Third, the LLM-as-a-Judge framework substantially enhances performance at early iterations (average correlation across tenors reaching 0.57), though subsequent iterations show degradation, emphasizing the need for human oversight in automated prompt optimization.
+First, even a naive zero-shot approach achieves substantial correlations (roughly 0.5) with actual market disagreement, particularly for medium and long-term maturities.
+This suggests LLMs can effectively decode central bank communication in economically meaningful ways.
+Second, the LLM-as-a-Judge framework substantially enhances performance at early iterations (average correlation across tenors reaching 0.57), though subsequent iterations show degradation, emphasizing the need for human oversight in automated prompt optimization.
 The stability of results across training and test datasets demonstrates that our approach captures fundamental communication-market relationships rather than memorizing patterns.
@@
 For market participants, it provides insights into the drivers of post-announcement volatility.
-For researchers, it opens the ”black box” of expectation formation by providing a micro-founded framework that explicitly models interpretive processes, moving beyond traditional reduced-form event studies.
-The methodology is readily adaptable to other central banks and policy contexts.
-Despite these contributions, a few limitations remain.
-The replicability of our results is partly constrained by the versioning and closed-source nature of contemporary LLMs, meaning that future iterations of the same model may yield slightly different outputs.
-A natural way forward would be to employ open-source models, which would allow for greater control, transparency, and long-term reproducibility.
-[^30] See Rogers and Luccioni, (2023) for a thorough discussion.
-Regarding potential contamination, while ECB press conferences are relatively niche compared to mainstream internet content, the possibility that fragments of such communication—or analyses thereof—were included in the model’s pretraining data cannot be completely dismissed.
-In our setup, we explicitly instruct the model to use only information available up to the date of each press conference; however, this constraint affects its responses, not its underlying parameters.
-Nevertheless, we consider the risk of substantial contamination to be limited, as discussions of post-ECB press conference volatility are relatively rare in the public domain, with only a small number of academic papers addressing the topic.
-Future research should focus on tackling these issues.
-Moreover, we consider these results a lower bound: testing with domain-specific LLMs or fine-tuned models could shed light on the trade-off between interpretive diversity and specialized expertise, potentially improving both the accuracy and the behavioral realism of simulated market reactions.
+For researchers, it opens the “black box” of expectation formation by providing a micro-founded framework that explicitly models interpretive processes, moving beyond traditional reduced-form event studies and rule-based models.
+The methodology is flexible and readily adaptable to other central banks and policy contexts.
+Despite these contributions, some limitations remain.
+The closed-source nature of leading LLMs constrains replicability, as future model versions may yield slightly different outputs.
+While we mitigate this problem through multi-model robustness testing, future research could explore open-source alternatives for greater transparency and reproducibility.
+Regarding potential contamination, while ECB press conferences are relatively niche compared to mainstream internet content, the possibility that fragments of such communication, or analyses thereof, were included in the model’s pretraining data cannot be completely dismissed.
+We explicitly instruct the model to use only information available up to each conference date; however, this constraint affects responses, not underlying parameters.
+Nevertheless, we consider substantial contamination risk limited given that discussions of post-ECB press conference volatility are rare in the public domain.
+Finally, domain-specific or fine-tuned models could potentially enhance performance by balancing interpretive diversity with specialized financial expertise, though this remains to be tested.
 In conclusion, this paper establishes a new paradigm for analyzing monetary policy communication at the intersection of artificial intelligence, behavioral finance, and central banking practice.
 By showing that LLMs can simulate heterogeneous trader interpretations and capture market disagreement patterns, we provide policymakers with a powerful tool to anticipate market reactions and refine communication strategies, while offering researchers a framework to better understand the micro-foundations of expectation formation.
-Appendix
+Appendix A: Additional Sensitivity Checks And Robustness Details
+6.1 Different Correlation Measures
+To ensure the robustness of our main findings, we conduct a supplementary analysis verifying that the observed relationship between synthetic and market-based disagreement is not dependent on the choice of the correlation metric.
+Our baseline analysis employs the Spearman rank correlation coefficient ($\rho$), which is well-suited for this context as it captures monotonic relationships and is robust to outliers and non-linearities.
+For a more comprehensive validation, we also consider two additional measures: the Pearson linear correlation coefficient ($r$) and Kendall’s rank correlation coefficient ($\tau$):
+  • Pearson’s $r$ quantifies the strength of a strictly linear relationship between two variables.
+Although more sensitive to outliers, it allows us to check whether the relationship could plausibly be explained within a linear framework.
+  • Kendall’s $\tau$ provides another non-parametric, rank-based alternative that evaluates concordance between two rankings.
+It is often preferred in smaller samples or in the presence of tied ranks, offering an additional robustness check alongside Spearman’s $\rho$.
+To assess the precision of these estimates, we compute 95% confidence intervals via a non-parametric bootstrap with 5,000 replications. Figure 14 presents the results.
+Figure 14: Comparison of Correlation Measures Between Synthetic and Market-Based Disagreement.
+Note: 95% confidence intervals computed via a non-parametric bootstrap with 5,000 replications.
+Across all three metrics, we find a statistically and economically significant positive association between our LLM-generated disagreement measure and market volatility.
+The 2-year tenor exhibits the strongest association in the rank-based measures (Spearman’s $\rho=0.53$; Kendall’s $\tau=0.37$), highlighting the simulation’s effectiveness at capturing uncertainty around the medium-term monetary policy path.
+Importantly, the bootstrapped 95% confidence intervals for all coefficients across all tenors comfortably exclude zero, underscoring statistical significance.
+While the magnitudes vary across correlation metrics, consistent with their differing assumptions, the overall conclusion remains unchanged.
+The alignment of results across linear and rank-based approaches demonstrates that the link we identify is a robust and stable feature of the data.
+6.2 Sensitivity to Market Volatility Window
+To ascertain whether our baseline findings are contingent upon the specific 3-day window used to measure market volatility, we conduct a robustness check using alternative 1-day and 2-day post-conference windows.
+The results shown in Figure 15 confirm that the positive and statistically significant relationship between our synthetic disagreement measure and market volatility is robust across all tested specifications.
+Figure 15: Correlation between Synthetic Disagreement and Market-Based Calculated Over Different Windows
+Note: Market- based definitions are equal to the min-max range in the day of the press conference, the average of the day of the conference and the following day, and within a three days period (baseline).
+95% confidence intervals computed via a non-parametric bootstrap with 5,000 replications.
+The analysis reveals two key patterns.
+First, for the 2-year tenor, the Spearman correlation is both high and remarkably stable (in the ballpark of 0.5).
+This stability reinforces the conclusion that the simulation is particularly adept at capturing uncertainty over the medium-term policy path, which appears to be the dominant factor in the days immediately following the communication event.
+Second, for the 10-year tenor, the correlation is strongest in the 1-day window (0.49) before declining slightly in the 2- and 3-day windows (0.44).
+This suggests our LLM-based framework is exceptionally effective at capturing the immediate, pure market interpretation of long-term policy signals.
+The subsequent modest decline in correlation may reflect the re-emergence of other long-term macroeconomic factors that influence the 10-year rate beyond the initial communication shock.
+Nevertheless, confidence intervals are largely overlapping.
+Consistent with our main findings, the 3-month tenor exhibits a statistically significant but weaker correlation across all windows.
+6.3 Model perturbation
+Across all three major language models tested, ChatGPT, Claude, and Gemini, the 2-year tenor synthetic value consistently achieves the highest correlations with actual market volatility, with Claude delivering the strongest performance at 0.586, followed by Gemini at 0.526 and ChatGPT at 0.440.
+This pattern reinforces our earlier findings that medium-term maturities, which primarily capture monetary policy expectations shaped by ECB forward guidance, are most amenable to LLM-based interpretation.
+The 10-year tenor shows more variable performance across models, with Claude again leading (0.563) but with a notably wider spread between the best and worst performing models (0.563 vs 0.287 for ChatGPT), suggesting that longer-term yield dynamics present greater challenges for consistent LLM interpretation.
+At the short end, the 3-month tenor exhibits the most uniform performance across models, with correlations ranging from 0.389 to 0.467, reflecting the inherent constraints of modeling near-term rates during periods of policy predictability.
+Importantly, Claude’s superior performance across all tenors may reflect architectural differences in processing complex financial text, though the substantial correlations achieved by all three models (consistently above 0.28) demonstrate that the core phenomenon, LLMs’ ability to capture interpretive heterogeneity in response to central bank communication, is not model-specific but rather represents a fundamental capability of this class of models.
+[^23] Initially, we had scraped some conferences included in the monetary policy section of the ECB website that were not actually about monetary policy (due to the ECB’s classification error).
+When running all the LLMs on those conferences, Claude was the only one that did not create the usual table, instead responding that those texts were not about monetary policy decisions.
+This demonstrates Claude’s superior content validation capabilities and suggests higher reliability in distinguishing relevant monetary policy communication from unrelated content.
+Figure 16: Individual Model Correlations with Market-Based Disagreement by Tenor
+Note: Spearman correlation coefficients between synthetic disagreement measures generated by individual LLM models (Gemini 2.5-Flash, ChatGPT 5-mini, and Claude Sonnet 4) and market-based post-conference OIS volatility across three tenors (3-month, 2-year, and 10-year).
+Each model generates synthetic disagreement by simulating 30 heterogeneous traders interpreting ECB press conference transcripts.
+Market-based disagreement is measured as the average min-max range of OIS rates in the 3 days following ECB press conferences.
+Sample covers 283 ECB Governing Council meetings from June 1998 to April 2025.
+6.4 Prompt perturbation
+Table 2: List of Prompt Variations
+Variation Type Key Changes Modified Section
+1 Minor Individual traders → distinct market participants Context
+2 Minor Trading decision → investment choice Context
+3 Minor Risk aversion order: High/Medium/Low → Low/Medium/High Trader Characteristics
+4 Minor Interpretation of conference → analysis of meeting Context
+5 Minor Markdown table → structured table format Guidelines
+6 Minor Confidence score → certainty level Task
+7 Minor Monetary policy decisions → interest rate policies Context
+8 Minor Maximize profit → optimize returns Context
+9 Minor Reordered behavioral biases (Anchoring first) Trader Characteristics
+10 Minor Press conference → policy meeting throughout Multiple sections
+11 Medium Formal academic language; market agents Context
+12 Medium Behavioral lenses; psychological profiles emphasis Context
+13 Medium Risk tolerance: Conservative/Moderate/Aggressive Trader Characteristics & Context
+14 Medium Policy signals; forward guidance; processing approach Context & Characteristics
+15 Medium Position-taking behavior; different output headers Task & Output
+Appendix B: LLM-as-Judge Implementation Details and Results
 Algorithm 1 LLM-as-a-Judge Prompt Optimization Algorithm for Trading Decisions

### v3

@@
 Abstract
-This paper develops a novel method to simulate financial market reactions to European Central Bank (ECB) press conferences using Large Language Models (LLMs).
-We create a behavioral, agent-based simulation of 30 synthetic traders, each with distinct risk preferences, cognitive biases, and interpretive styles.
-These agents interpret ECB communication and forecast Euro interest rate swap levels at 3-month, 2-year, and 10-year maturities, with cross-sectional variation serving as a measure of market uncertainty.
-Using a comprehensive dataset of 283 ECB press conferences (1998-2025), we evaluate three prompting strategies: naive zero-shot, few-shot (enriched with historical context), and an advanced iterative ”LLM-as-a-Judge” framework.
-We find that even the naive approach achieves substantial correlations (roughly 0.5) between synthetic disagreement and actual post-announcement OIS volatility, particularly for medium- and long-term maturities.
-The LLM-as-a-Judge framework further enhances performance upon initial application, reaching correlations of almost 0.6 and identifying key drivers of volatility.
-Our results demonstrate that LLM-based agents can effectively capture the heterogeneous interpretive processes through which monetary policy signals propagate to financial markets, moving beyond traditional reduced-form event studies.
-This provides central banks with a practical tool to ex-ante evaluate communication strategies and anticipate market reactions, while offering researchers a micro-founded framework for understanding expectation formation in response to central bank communication.
+Traditional high-frequency identification of monetary policy communication effects operates ex-post, precluding evaluation of alternative strategies before publication.
+In this paper, we introduce an ex-ante framework that simulates heterogeneous market reactions to central bank communication before release.
+Our methodology employs Large Language Models (LLMs) to construct an agent-based simulation of 30 synthetic traders with heterogeneous risk preferences, cognitive biases, and interpretive styles.
+These agents process European Central Bank (ECB) press conference transcripts and forecast Euro interest rate swap rates across 3-month, 2-year, and 10-year maturities.
+Cross-sectional forecast dispersion provides a model-based measure of market disagreement, validated against realized overnight index swap (OIS) volatility.
+Analyzing 283 ECB press conferences (June 1998-April 2025), we document Spearman correlations of approximately 0.5 between simulated and realized disagreement, rising to 0.6 under iterative prompt optimization.
+Results prove robust across prompting strategies, are temporally stable across training and holdout samples, and fare significantly better than simple language complexity scores.
+For central banks, the framework provides an operational tool to anticipate communication-induced volatility before release, thus enabling ex-ante language refinement.
+For researchers, it offers a micro-founded alternative to reduced-form event studies, explicitly modeling the heterogeneous interpretive processes through which policy signals are transmitted to asset prices.
 Keywords: large language models, monetary policy surprises, monetary policy communication
@@
 […] We thoroughly expect to be made largely redundant by computers eventually, but perhaps not for a few years to come.” Romer and Romer, (2023)
-Recent advancements in Large Language Models (LLMs) have opened up a plethora of possibilities for economic and financial research.
-Due to the way these models are trained, on vast and diverse textual data, and their ability to generate contextually rich and realistic responses, LLMs can now be viewed as implicit computational representations of human behavior.
-Much like the economic construct of homo economicus, these models can be endowed with information, placed in hypothetical scenarios, and observed for their responses (Horton,, 2023; Kazinnik and Sinclair,, 2025).
-Despite their growing capabilities, the application of LLMs in finance has been relatively limited so far.
-[^1] A notable exception is, for example, Lopez-Lira and Tang, (2023).
-Yet, the financial domain may be one of the most promising areas for their use.
-Unlike traditional economic models that assume rational agents and efficient markets, real-world financial markets are often driven by sentiment, narrative, and behavioral biases, factors that are difficult to capture with purely mathematical models, especially in high-frequency settings.
-[^2] Agent-based models (ABMs) have emerged as a flexible framework to incorporate such heterogeneity (Farmer and Foley,, 2009).
-In the context of monetary policy, ABMs have been used to simulate expectation formation, market volatility, and policy transmission (Delli Gatti et al.,, 2011).
-However, most ABMs rely on stylized behavioral rules imposed by the researcher (Horton,, 2023).
-A key area where LLMs can offer significant value is in modeling reactions to central bank communication, a cornerstone of modern monetary policy.
-It is well-established that the effectiveness of monetary policy hinges not just on the decisions themselves, but on the language used and its interpretation by economic agents.
-Press conferences, speeches, and official statements play a pivotal role in shaping market expectations and, consequently, asset prices (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019).
-Previous research, has moved beyond the general direction of the market response to highlight the crucial role of response heterogeneity; in the Euro area, high disagreement following a statement transmits to financial markets through a separate channel compared to the standard monetary policy surprise channel, and can dampen the intended impact of policy tightening by lowering the term premium of the yield curve (Collodel and Kunzmann,, 2025).
-[^3] Bauer et al., (2021) find similar results for the US.
-Capturing this nuanced, belief-driven heterogeneity is a challenge for traditional models but is a natural fit for the capabilities of LLMs.
-This paper introduces a novel approach to modeling market reactions to ECB communication by leveraging the capabilities of LLMs.
-We develop a behavioral, agent-based simulation of 30 synthetic traders through Gemini LLM, one of Google’s leading AI products.
-These agents respond to the texts of ECB press conferences by forecasting Euro interest rate swap levels across three key maturities, short (3m), medium (2y), and long-term (10y).
-The main idea is that the model can capture a range of possible interpretations of these conferences.
-Thus, the variation in the resulting forecasts, measured by their standard deviation, serves as a proxy for market uncertainty or disagreement.
-By correlating this synthetic measure with actual market outcomes, we study the ability of LLMs to replicate the heterogeneity of investors’ beliefs.
-To further explore how the design of prompting influences this correlation, we implement three distinct prompting strategies.
-The first is a naive or “zero-shot” prompt (Radford et al.,, 2019), which serves as the baseline: each agent is assigned a simple, pre-defined level of risk aversion, a set of cognitive biases, and a unique interpretive style.
-The second is a “few-shot” prompt (Brown et al.,, 2020), which enriches the agent’s context by including the past three ECB press conferences along with the corresponding changes in market-based disagreement resulting from them.
-Finally, we introduce an LLM-as-a-Judge prompt, our most advanced strategy (Zheng et al.,, 2023).
-In this iterative setup, a separate LLM, the Judge, acts as a critical evaluator, systematically reviewing the analyst prompt and its resulting in-sample correlation score.
-The Judge then rewrites the prompt to improve alignment with market reactions.
-This revised prompt is fed back into the Analyst LLM, forming a feedback loop that continues over different iterations.
-The different prompt versions are subsequently evaluated out-of-sample to assess their robustness and decrease concerns relative to model over-fitting.
-We find that LLM-based agents can effectively replicate the complex and varied reactions of investors to ECB press conferences, achieving a remarkable correlation of around 0.5 between the synthetic disagreement measure and actual market outcomes for medium- and long-term maturities, a substantial result given the inherent complexity and noise of financial markets.
-Contrary to expectations, a more detailed few-shot prompt does not improve this correlation; however, the LLM-as-a-Judge framework significantly enhances predictive accuracy upon its initial application, with performance degrading thereafter.
-These results suggest that LLM-based agent simulations may offer policymakers a useful tool for evaluating central bank communications and identifying language that could generate market disagreement prior to public release.
-By moving beyond traditional sentiment analysis, our approach contributes to understanding how monetary policy signals are perceived and transmitted through financial markets.
-The framework integrates insights from macroeconomic communication, behavioral finance, and artificial intelligence to help scholars and policymakers better understand the financial market impact of monetary policy announcements.
+Central banks face a fundamental communication dilemma: they cannot test how markets will interpret their statements before release.
+Upon release, policy language triggers immediate and irreversible market reactions.
+Traditional high-frequency identification methods measure these reactions ex-post, i.e., revealing what happened (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019), but offering no guidance before release.
+This constraint is costly.
+Ambiguous communication can amplify market volatility, hindering financial stability and undermining policy transmission, yet policymakers lack tools to anticipate these effects during the drafting process when language remains modifiable.
+This paper introduces the first ex-ante framework for simulating heterogeneous market reactions to monetary policy communication.
+We employ Large Language Models (LLMs) to construct an agent-based simulation of 30 synthetic traders, each endowed with distinct risk preferences, cognitive biases, and interpretive styles.
+These agents process European Central Bank (ECB) press conference transcripts and forecast Euro interest rate swap rates across three key maturities representing the whole yield curve spectrum (3-month, 2-year, 10-year).
+Cross-sectional forecast dispersion provides a model-based measure of market disagreement, which we validate against realized overnight index swap (OIS) volatility across 283 ECB communications spanning the period 1998-2025.
+Our approach addresses a critical gap in the monetary policy communication literature.
+While extensive research documents that central bank language moves markets (Gurkaynak,, 2005; Altavilla et al.,, 2019), and that disagreement about policy stance affects transmission through distinct channels, including term premium compression (Tillmann,, 2020; Bauer et al.,, 2021; Collodel and Kunzmann,, 2025), existing methods remain fundamentally retrospective.
+They identify communication shocks from realized price movements, leaving the interpretive process itself as a black box.
+We attempt to open this box by explicitly modeling how heterogeneous agents process policy language, generating predictions of disagreement before markets react.
+The framework builds on recent research demonstrating that LLMs, being trained on extensive textual corpora, can effectively simulate human interpretive behavior (Horton,, 2023; Kazinnik and Sinclair,, 2025).
+Unlike traditional agent-based models, which rely on stylized behavioral rules (Delli Gatti et al.,, 2011), LLM-based agents can directly process natural language, extracting meaning from the same textual signals that influence market participants.
+This ability is especially valuable in the context of monetary policy, where communication effects hinge on narrative and linguistic nuance rather than purely mathematical relationships.
+Financial markets are shaped by sentiment, biases, and interpretive style heterogeneity, features that are difficult to capture with conventional models but align naturally with the strengths of LLMs.
+We implement three prompting strategies of increasing sophistication.
+The baseline “zero-shot” approach assigns to each agent fixed characteristics (risk aversion, cognitive biases, interpretive style) and evaluates their raw reactions to press conferences, without historical context.
+The “few-shot” specification enriches agents with the prior three conferences and associated market reactions, testing whether historical calibration improves prediction.
+Our most advanced approach employs an “LLM-as-a-Judge” framework (Zheng et al.,, 2023): a separate LLM iteratively critiques and rewrites the analyst prompt based on in-sample correlation with the market-based measure, then validates optimized prompts out-of-sample to assess overfitting.
+We find that even the baseline specification achieves Spearman correlations of approximately 0.5 between simulated and realized disagreement for medium- and long-term maturities.
+Performance peaks at the 2-year tenor, consistent with this maturity’s elevated sensitivity to forward guidance, and rises to 0.6 under initial LLM-as-a-Judge optimization.
+Contrary to expectations, few-shot prompting does not improve performance.
+Results prove robust across model architectures.
+prompt choices, and are temporally stable in holdout samples.
+The framework contributes to policy practice and economic research.
+For central banks, it provides an operational tool to anticipate communication-induced volatility before release, enabling ex-ante language refinement and counterfactual testing of alternative phrasings.
+For researchers, it offers a micro-founded methodology for studying expectation formation, moving beyond reduced-form event studies to explicitly model the heterogeneous belief processes through which policy signals transmit to asset prices.
+By demonstrating that interpretive disagreement is both quantifiable and predictable from text alone, we establish that communication strategy can be optimized systematically rather than refined only through costly trial and error.
 The remainder of the paper proceeds as follows.
-Section 2 reviews the relevant literature on monetary policy communication, behavioral expectations, and LLM applications in economics.
-Section 3 outlines our methodology, including data collection of ECB press conferences and market reactions, construction of synthetic agent personas, and our three prompting strategies.
-Section 4 presents our empirical results, analyzing agent forecast dispersion patterns, correlation with market volatility measures, and comparative performance across prompting approaches.
-Section 5 probes the robustness of our results to model and prompt variation.
-Lastly, Section 6 concludes with policy implications for central bank communication strategies and directions for future research.
+Section 2 reviews relevant literature on monetary policy communication, behavioral expectations, and LLM applications in economics.
+Section 3 outlines our methodology, including the construction of synthetic agents and prompting strategies.
+Section 4 presents empirical results on forecast dispersion patterns and correlation with market volatility.
+Section 5 examines robustness to model and prompt variation.
+Section 6 concludes with implications for central bank communication design and future research directions.
 2 Literature Review
@@
 In these studies, changes in risk-free market rates measured before and after announcements serve as proxies for monetary policy shocks.
+[^1] For the sake of precision, the authors decompose these changes into orthogonal factors, before treating them as shocks.
 In the euro area, similar evidence shows that ECB press conferences contain significant informational content beyond the policy decision itself (Altavilla et al.,, 2019).
@@
 Yet two key limitations remain.
-First, behavioral rules are typically imposed ex ante as stylized heuristics (Horton,, 2023), rather than learned or emergent.
+First, behavioral rules are typically imposed ex ante as stylized heuristics, rather than learned or emergent (Horton,, 2023).
 Second, conventional ABMs cannot interpret qualitative information directly, despite the fact that monetary policy signals are largely communicated in natural language.
@@
 We focus on ECB press conferences, which constitute the primary vehicle for conveying monetary policy decisions in the Euro area.
-Following the literature on high-frequency identification of monetary policy shocks (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019), we define event dates as the days of regularly scheduled ECB Governing Council monetary policy meetings, limiting the sample to those that include both a written policy statement and a subsequent press conference by the ECB President.
+Following the literature on high-frequency identification of monetary policy shocks (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019), we define event dates as the days of regularly scheduled ECB Governing Council monetary policy meetings.
 For each event, we collect the full transcript of the press conference as published on the ECB website as well as the associated OIS post-conference volatility for the short (3-months), medium (2-years), and long (10-years) term.
@@
 Zero-Shot Prompt:
-The baseline approach, where the agent receives only the ECB press conference transcript along with its fixed attributes.
+The baseline approach, where the agent receives only the ECB press conference transcript.
 This approach tests whether LLMs possess intrinsic capabilities to decode monetary policy communication without explicit calibration.
@@
 Augments the baseline by including the transcripts of the three most recent ECB press conferences and the associated historical changes in market-based disagreement, providing a short memory of past communication–market interactions.This design tests whether contextual examples improve forecasting accuracy, following established few-shot learning principles.
-[^8] Few-shot prompting has demonstrated superior performance over zero-shot approaches across multiple economic forecasting applications, particularly when fine-tuning is computationally infeasible.
+[^6] Few-shot prompting has demonstrated superior performance over zero-shot approaches across multiple economic forecasting applications, particularly when fine-tuning is computationally infeasible (Hansen et al.,, 2024).
 LLM-as-a-Judge Prompt:
@@
 We examine the zero-shot baseline (Section 4.1), the few-shot approach with historical context (Section 4.2), and the iterative LLM-as-a-Judge framework (Section 4.3), focusing on correlations between synthetic disagreement and actual market volatility across different yield curve tenors.
-We conclude with a comparison against simple textual benchmarks (Section LABEL:sec:complexity) to demonstrate the value of modeling heterogeneous agent interpretations compared to less sophisticated approaches.
 4.1 Zero-Shot Prompt
@@
 Forecast directions are aggregated across agents.
-The simulation uses a sample of 283 ECB press conference (June 1998-April 2025) transcripts and prompt 2.
+The simulation uses a sample of 283 ECB press conference (June 1998-April 2025) transcripts and the baseline prompt (Figure 2).
 Panel 4 plots the share of agents predicting an increase, decrease, or no change in rates after each ECB press conference.
@@
 Dispersion is measured as the cross-sectional standard deviation of rate predictions.
-The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025) and prompt 2.
+The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025) and the baseline prompt (Figure 2).
 Highlighted points represent observations exceeding the 95th percentile of dispersion for each tenor.
-Panel 5 shows the cross-sectional standard deviation of forecasts, revealing distinct phases that align closely with monetary policy regimes and market stress episodes.
+Figure 5 shows the cross-sectional standard deviation of forecasts, revealing distinct phases that align closely with monetary policy regimes and market stress episodes.
 Dispersion contracts systematically during periods of unambiguous policy stance, such as the post-crisis easing phase and much of the zero lower bound (ZLB) era, when forward guidance provided clear anchoring.
@@
 Note: Correlation between the output of an LLM model prompted in a zero-shot setting using Gemini 2.5-Flash, with a temperature parameter of 1, and a market-based measure of monetary policy uncertainty from Collodel and Kunzmann, (2025).
-The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025) and prompt 2.
+The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025) and the baseline prompt (Figure 2).
 Rolling correlation is calculated using a 12-month window.
@@
 This few-shot approach aims to help agents anchor their interpretations in observed patterns between ECB communication and market reactions, following evidence that historical examples can improve LLM forecasting performance in other economic contexts (Hansen et al.,, 2024).
-The few-shot specification does not materially improve correlation with market volatility compared to the naive model: correlations remain essentially unchanged across maturities, 2-year (0.52 vs.
-0.48), 10-year (0.45 vs.
-0.46), and 3-month (0.38 vs.
+The few-shot specification does not materially improve correlation with market volatility compared to the naive model: correlations remain essentially unchanged across maturities, 2-year (0.53 vs.
+0.48), 10-year (0.44 vs.
+0.46), and 3-month (0.39 vs.
 0.38), and none of these differences are statistically significant.
@@
 Dispersion is measured as the cross-sectional standard deviation of rate predictions.
-The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025), prompt 2 and provides the text of the three previous conferences as well as the ensuing market-based standard deviation of OIS rates.
+The simulation uses a sample of 283 ECB press conferences (June 1998-April 2025), the baseline prompt (Figure 2), and provides the text of the three previous conferences as well as the ensuing market-based standard deviation of OIS rates.
 Highlighted points represent observations exceeding the 95th percentile of dispersion for each tenor.
@@
 Figure 8: Correlation with market-based measures across iterations - training and test set
-Note: Each point reflects the average Spearman correlation across tenors (3months, 2-years and 10-years) between LLM-generated rate Euro OIS forecasts and market-based monetary policy uncertainty from Collodel and Kunzmann, (2025) after ECB GovC press conferences.The x-axis indicates the number of prompt iteration in a LLM-as-Judge framework.
+Note: Each point reflects the average Spearman correlation across tenors (3months, 2-years and 10-years) between LLM-generated standard deviation of Euro OIS rate forecasts and market-based monetary policy uncertainty from Collodel and Kunzmann, (2025) after ECB GovC press conferences.The x-axis indicates the number of prompt iteration in a LLM-as-Judge framework.
 The simulation uses Gemini 2.5-Flash with a temperature of 1, and incorporates a feedback loop via a separate LLM (“Judge”) that uses Gemini 2.5-Pro.
@@
 Note: Time series comparison of LLM-predicted forecast dispersion and actual market volatility following ECB press conferences, segmented by prompt iteration and tenor.
-Simulations employ Gemini 2.5-Flash with temperature = 1 and a Judge-based feedback loop that iteratively refines prompts over 5 iterations.
+Simulations employ Gemini 2.5-Flash with temperature = 1 and the second prompt generated by a Judge-based feedback loop that iteratively refines prompts over 5 iterations.
 Forecast dispersion represents the standard deviation across 5 simulation runs per prompt per conference.
@@
 Second, we assess model architecture stability across different LLM systems (Gemini, ChatGPT, and Claude) to ensure our results reflect systematic patterns captured by this class of models rather than artifacts of a particular model’s design.
-Third, we examine prompt specification stability, addressing the well-documented fickleness of LLMs to minor wording changes.
+Third, we examine prompt specification stability, addressing the well-documented fickleness of LLMs to minor wording changes (Loya et al.,, 2023; Sclar et al.,, 2024).
 5.1 Simple Benchmark Comparison
@@
 The Flesch-Kincaid score is a standard measure of text readability that estimates the U.S.
-school grade level required to understand a given text, with higher scores indicating more complex language.
+school grade level required to understand a given text, with higher scores indicating more complex language and longer, more technical sentences.
 In related work, Mumtaz et al., (2023) show that communication complexity at the Bank of England is associated with higher gilt yield volatility, with complex statements producing significantly larger increases in the standard deviation of returns compared to simpler ones.
 Our analysis tests whether this relationship holds for ECB communications and how it compares to our LLM-based approach.
-Figure 10: Correlation between Flesh-Kincaid score and post ECB conference OIS volatility
-Note: Volatility in percentage points.
-The correlations between Flesch-Kincaid scores and post-conference OIS volatility are 0.15, 0.26, and 0.06 for the 3-month, 2-year, and 10-year tenors, respectively (Figure 10).
-These results suggest that readability and complexity might be part of the story: more complex communication can amplify market uncertainty.
-Nevertheless, this dimension captures only one small facet of the broader interpretive process.
-By construction, readability indices cannot capture policy content, such as the magnitude of the monetary policy surprise, or strategy shifts, like the abandonment of forward guidance, so their explanatory power remains limited.
-For example, during the Covid-19 pandemic, policy statements became highly complex owing to the unprecedented nature of the health shock, yet forward guidance kept expectations well anchored and first-moment surprises remained small, thereby limiting volatility.
-In contrast, our LLM-based framework delivers substantially higher correlations by capturing the heterogeneous interpretive processes through which market participants react to central bank communication.
+Figure 10: Flesch-Kincaid complexity and post-ECB conference OIS volatility
+Note: Market volatility measured as the average min-max range of OIS rates in the 3 days following ECB press conferences.
+Outliers (greater than 99th percentile) excluded.
+LOESS smoothing reflects rank-based relationship.
+Spearman correlation ($\rho$) shown in the top-right corner of each panel.
+Figure 10 presents the relationship between Flesch-Kincaid complexity scores and market-based disagreement across maturities.
+The Spearman correlations are modest: 0.23 for both 3-month and 2-year tenors, and essentially zero (0.03) for the 10-year tenor.
+These results suggest that while syntactic complexity may play a role, it explains only a small fraction of market disagreement.
+Moreover, the relationship appears non-linear, particularly at the 2-year maturity, where disagreement peaks at moderate complexity levels before declining, a pattern inconsistent with a simple ”more complexity equals more uncertainty” narrative.
+This weak correlation between readability metrics and market outcomes highlights a fundamental limitation: by construction, readability indices measure only surface-level linguistic features, such as sentence length, word difficulty, and grammatical structure, but cannot capture the substantive dimensions that drive interpretive disagreement i.e.
+policy content (the magnitude of monetary surprises), strategic shifts (such as the abandonment of forward guidance), internal contradictions between prepared statements and Q&A sessions, or the excessive use of hedging language.
+A paramount example is the COVID-19 pandemic: during this period, policy statements became highly complex due to the unprecedented nature of the health shock; yet, strong forward guidance kept expectations well anchored, and first-moment surprises remained small, thereby limiting volatility despite high Flesch-Kincaid scores.
+In stark contrast, our LLM-based framework achieves correlations of approximately 0.5 for medium- and long-term tenors, more than double the readability-based benchmark.
+This substantial improvement reflects the LLM’s ability to capture what readability metrics cannot: how heterogeneous agents actually interpret policy signals, decode ambiguous forward guidance, and react to conflicting messages within the same communication.
+These findings validate our core methodological contribution: understanding market reactions to central bank communication requires modeling the interpretive processes of market participants, not merely measuring syntactic complexity.
 5.2 Model Architecture Robustness
@@
 6 Conclusions
-This paper introduces a novel methodology using Large Language Models to simulate heterogeneous market reactions to ECB press conferences.
-Through an agent-based framework with 30 synthetic traders, each endowed with distinct behavioral characteristics, we demonstrate that LLMs can effectively capture market disagreement patterns in response to monetary policy communication.
-Our key findings advance both theoretical understanding and practical application in monetary economics.
-First, even a naive zero-shot approach achieves substantial correlations (roughly 0.5) with actual market disagreement, particularly for medium and long-term maturities.
-This suggests LLMs can effectively decode central bank communication in economically meaningful ways.
-Second, the LLM-as-a-Judge framework substantially enhances performance at early iterations (average correlation across tenors reaching 0.57), though subsequent iterations show degradation, emphasizing the need for human oversight in automated prompt optimization.
-The stability of results across training and test datasets demonstrates that our approach captures fundamental communication-market relationships rather than memorizing patterns.
-These results have significant practical implications.
-For central banks, our framework offers a powerful ex-ante tool to evaluate communication strategies and identify potentially market-disrupting language before release.
-For market participants, it provides insights into the drivers of post-announcement volatility.
-For researchers, it opens the “black box” of expectation formation by providing a micro-founded framework that explicitly models interpretive processes, moving beyond traditional reduced-form event studies and rule-based models.
-The methodology is flexible and readily adaptable to other central banks and policy contexts.
-Despite these contributions, some limitations remain.
-The closed-source nature of leading LLMs constrains replicability, as future model versions may yield slightly different outputs.
-While we mitigate this problem through multi-model robustness testing, future research could explore open-source alternatives for greater transparency and reproducibility.
-Regarding potential contamination, while ECB press conferences are relatively niche compared to mainstream internet content, the possibility that fragments of such communication, or analyses thereof, were included in the model’s pretraining data cannot be completely dismissed.
-We explicitly instruct the model to use only information available up to each conference date; however, this constraint affects responses, not underlying parameters.
-Nevertheless, we consider substantial contamination risk limited given that discussions of post-ECB press conference volatility are rare in the public domain.
-Finally, domain-specific or fine-tuned models could potentially enhance performance by balancing interpretive diversity with specialized financial expertise, though this remains to be tested.
-In conclusion, this paper establishes a new paradigm for analyzing monetary policy communication at the intersection of artificial intelligence, behavioral finance, and central banking practice.
-By showing that LLMs can simulate heterogeneous trader interpretations and capture market disagreement patterns, we provide policymakers with a powerful tool to anticipate market reactions and refine communication strategies, while offering researchers a framework to better understand the micro-foundations of expectation formation.
+Central banks face a fundamental communication challenge: they cannot test how markets will interpret their statements before release.
+Traditional high-frequency identification methods measure market reactions ex-post, revealing what happened but offering no guidance for ex-ante communication design.
+This constraint is particularly costly given mounting evidence that interpretive disagreement affects financial stability and monetary transmission.
+This paper addresses the gap by developing the first ex-ante framework for simulating heterogeneous market reactions to central bank communication using Large Language Models.
+We construct an agent-based simulation of 30 synthetic traders with distinct behavioral characteristics who process ECB press conference transcripts and forecast Euro interest rate swap rates across the yield curve.
+Cross-sectional forecast dispersion provides a model-based measure of market disagreement that can be computed before any communication event occurs.
+Our empirical analysis of 283 ECB press conferences (June 1998–April 2025) yields three main findings.
+First, even without historical calibration, the zero-shot specification achieves Spearman correlations of approximately 0.5 with realized OIS volatility, demonstrating that LLMs possess intrinsic capabilities to decode central bank communication in economically meaningful ways.
+Second, the relationship proves strongest for the 2-year tenor ($\rho$ = 0.53), consistent with this maturity’s elevated sensitivity to forward guidance that dominated ECB communication for much of the period under review.
+The hierarchy across maturities, with medium-term rates most predictable, followed by long-term, then short-term, reflects fundamental differences in how policy signals transmit across the yield curve.
+Third, an adaptive “LLM-as-a-Judge” framework raises correlations to 0.6 at early iterations, though subsequent refinements degrade performance, demonstrating both the potential and limits of automated prompt optimization without human oversight.
+These findings carry direct policy implications.
+For central banks, the framework provides an operational tool to anticipate communication-induced volatility before release, enabling systematic evaluation of alternative phrasings, identification of potentially ambiguous language, and quantitative assessment of how textual changes would affect expected market disagreement (refer to section Appendix C: Counterfactual Communication Analysis in Appendix C for a simple counterfactual analysis).
+This transforms the communication strategy from reactive refinement, based on costly market reactions, to proactive optimization during the drafting process.
+For market participants, the methodology offers insights into drivers of post-announcement volatility and a framework for evaluating their own interpretive biases.
+For researchers, it establishes a micro-founded approach to studying expectation formation that explicitly models heterogeneous belief processes, moving beyond reduced-form event studies.
+The methodology’s adaptability extends its relevance beyond the ECB.
+The framework transfers naturally to other major central banks, Federal Reserve, Bank of England, Bank of Japan, enabling comparative analysis of how communication styles, institutional frameworks, and forward guidance regimes affect interpretive disagreement across monetary policy contexts.
+The broader implications extend to how we conceptualize the role of AI in economic research and policy.
+Our results demonstrate that LLMs can serve not merely as statistical pattern recognizers but as computational representations of human interpretive processes, in line with Horton, (2023).
+Nevertheless, important limitations warrant acknowledgment and suggest directions for future research.
+First, replicability faces constraints from LLM versioning and closed-source architectures, though our cross-model robustness checks and public release of prompts partially mitigate these concerns.
+Future work should prioritize open-source implementations for full transparency and reproducibility.
+Second, while we cannot completely rule out training data contamination, fragments of ECB communication or related analyses may have appeared in pretraining corpora, the rarity of detailed discussions about post-ECB volatility in public sources, combined with our temporal instructions limiting the model to information available at each conference date, substantially reduces this risk.
+Third, our reliance on general-purpose models may understate achievable performance.
+Domain-specific fine-tuning on central bank and financial text, or development of specialized architectures optimized for monetary policy interpretation, could significantly enhance accuracy.
+In conclusion, this paper establishes that LLM-based agent simulations can quantitatively reproduce market disagreement patterns arising from central bank communication with economically significant accuracy.
+By demonstrating that interpretive heterogeneity is both measurable and predictable from text alone, we provide central banks with a practical tool for ex-ante communication evaluation and offer researchers a framework for studying the micro-foundations of expectation formation in monetary policy transmission.
 Appendix A: Additional Sensitivity Checks And Robustness Details
@@
 Table 3: Correlation by prompt iteration (continued)
+Appendix C: Counterfactual Communication Analysis
+6.5 Motivation and Design
+Having established that LLM-based agents can provide central banks with an ex-ante assessment of the disagreement their communication is likely to generate, we now turn to a complementary question: Can central banks leverage LLMs to systematically refine communication drafts and thereby reduce ensuing interpretive uncertainty?
+To answer this question, we design a simple counterfactual experiment.
+The experiment proceeds in three stages.
+First, we identify 30 high-volatility press conferences from our sample, defined as those with the highest average post-announcement OIS volatility.
+These high-volatility episodes are precisely the communications that matter the most for central banks, making them the ideal sample for evaluating whether linguistic clarity through LLM refinement can reduce market uncertainty.
+Second, we employ ChatGPT 5o-mini to edit the text of these conferences, with explicit instructions to reduce ambiguity while maintaining the substantive policy message (Figure 18).
+Third, we submit both the original and edited transcripts to our Gemini 2.5-Flash simulation framework, generating synthetic disagreement measures for each version.
+The change in disagreement between the two versions, measured both in percentage terms and absolute basis points, serves as our Average Treatment Effect (ATE).
+Figure 18: ChatGPT 5o-mini Editing Prompt
+Counterfactual Editing Prompt You are editing an ECB press conference to reduce ONLY linguistic ambiguity.
+STRICT CONSTRAINTS - DO NOT CHANGE: - Any policy decisions (interest rates, asset purchases, programs) - Any economic data, statistics, or forecasts - The overall hawkish/dovish stance - Substantive content or meaning ONLY EDIT: 1.
+**Vague temporal phrases**: sufficiently long duration -> Use context from same text to infer timeframe some time’, appropriate period -> Make more specific using surrounding discussion 2.
+**Internal contradictions**: If prepared statement and Q&A conflict -> Align using the clearer version 3.
+**Excessive hedging**: Multiple qualifiers (’may potentially consider’) -> Simplify if context supports IMPORTANT: Base any specificity ONLY on information already in the text.
+Do NOT invent dates, numbers, or commitments.
+ORIGINAL TEXT ({date}): {text} EDITED VERSION:
+Note: Prompt used to generate counterfactual press conference transcripts with reduced linguistic ambiguity.
+The editor (ChatGPT-5o) is constrained to preserve all substantive policy content while only clarifying vague language, resolving contradictions, and reducing excessive hedging.
+Placeholders {date} and {text} are replaced with the actual conference date and transcript for each editing task.
+This cross-LLM design is the pillar of our counterfactual analysis.
+By using different models for editing (ChatGPT 5o-mini) and evaluation (Gemini 2.5-Flash), we mitigate a critical concern: that the evaluator might mechanically align with the editor’s interpretation simply because both tasks are performed by the same model.
+In simple words, if we used a single LLM for both editing and evaluation, we would not be able to distinguish genuine clarity improvements from the model’s tendency to find its own edits more coherent, a form of confirmation bias at the algorithmic level to which LLMs are prone.
+The cross-model design ensures that disagreement reduction reflects genuine improvements in textual clarity rather than model-specific interpretive consistency.
+6.6 Results
+Figure 19 shows the ATEs. Panel (a) shows the mean disagreement change in basis points, while panel (b) displays the percentage change across the three OIS tenors.
+Figure 19: Mean Disagreement Change by Tenor
+(a) Basis Points Reduction
+(b) Percentage Reduction
+Note: Mean disagreement change following linguistic edits by ChatGPT, evaluated by Gemini 2.5-Flash across three OIS tenors (3-month, 2-year, and 10-year).
+Panel (a) shows the reduction in basis points, while panel (b) shows the percentage reduction in forecast disagreement (reduction = positive value).
+Error bars represent 95% confidence intervals from one-sided t-tests.
+Disagreement is measured as the cross-sectional standard deviation of rate forecasts across 30 simulated heterogeneous traders interpreting ECB press conference transcripts.
+Sample covers 30 highest-volatility ECB Governing Council meetings in the period 1998-2025.
+The editing intervention achieves statistically significant reductions in disagreement at the 2Y and 10Y maturity, while the 3M tenor uncertainty slightly diminishes, but the impact is highly volatile across conferences.
+For the 10-year OIS, editing reduces disagreement by an average of 7.6 basis points (27.8% of the original volatility), while the 2-year tenor exhibits a reduction of 5.9 basis points (21.4%).
+The 3-month tenor shows a more modest, yet still meaningful, reduction of 2.1 basis points (4.44%).
+The maturity structure of these effects aligns with previous results.
+Longer-term rates are more sensitive to forward guidance and economic outlook assessments, which are precisely the textual elements most amenable to clarification.
+Short-term rates, by contrast, are more tightly anchored to the immediate policy decision and exhibit less scope for interpretive disagreement even in the baseline scenario.
+Figure 20 visualizes the conference-by-conference pattern of disagreement change.
+The heatmap clearly identifies specific press conferences where editing achieved substantial reductions (shown in dark blue) versus those where the impact was minimal or even counterproductive (shown in dark red).
+Figure 20: Change in Disagreement by Press Conference
+Note: Percentage reduction in disagreement for each of the 30 conferences with the highest market-based disagreement, by tenor (3-month, 2-year, and 10-year).
+ChatGPT (5o-mini) edits the original ECB press conference transcripts to reduce linguistic ambiguity, then Gemini (2.5-Flash) agents evaluate both original and edited versions.
+Disagreement is measured as the cross-sectional standard deviation of rate forecasts across 30 heterogeneous agents.
+Darker colors indicate larger changes in disagreement.
+Negative values indicate cases where editing increased disagreement.
+Two patterns emerge from the heatmap.
+First, the predominance of blue tiles indicates that editing reduces disagreement in the vast majority of conferences across all tenors.
+Second, when editing fails to reduce disagreement, or even increases it, these instances tend to occur consistently across all three maturities rather than being tenor-specific.
+The latter cases cluster in three distinct periods: the GFC (late 2008-early 2009), the recent tightening cycle (December 2022), and the early ECB period (late 1998), suggesting that LLM editing is less effective during episodes of extreme market stress or high institutional uncertainty.
+This result is consistent with a simple theoretical model of expectation formation under heterogeneous information processing.
+In the aforementioned setting, editing primarily reduces textual ambiguity by clarifying hedging language and resolving contradictions.
+Under normal conditions, this effectively compresses disagreement as agents converge toward a clearer public signal.
+However, during crisis episodes, agents enter the press conference with wildly divergent priors about the economic state and appropriate policy stance: when prior dispersion is so large, even clear communication cannot eradicate disagreement.
+The synchronous failures across all tenors confirm that this prior dispersion affects expectations uniformly across the yield curve, rather than reflecting tenor-specific uncertainty.
+Several considerations inform the interpretation of these results.
+First, the editing protocol we employ is necessarily stylized.
+Real-world communication involves trade-offs between clarity and flexibility, and central banks may deliberately preserve ambiguity to avoid premature commitment or to accommodate internal disagreement among policymakers.
+Our simplified editing instructions abstract from these strategic considerations.
+Second, human evaluation rather than LLM based appraisal would constitute the gold standard for validating our findings.
+While our LLM-based measurement provides scalable and replicable disagreement estimates, we cannot definitively rule out that both the editor (ChatGPT) and evaluator (Gemini) share systematic biases in how they process certain linguistic features.
+A robust extension would involve having professional economists or central bank communication specialists assess: (i) whether the edited transcripts preserve the intended policy message, (ii) whether the edits are institutionally and legally permissible, and (iii) whether human readers perceive the edited versions as genuinely clearer.
+Such validation would confirm whether the disagreement reductions we observe reflect actual improvements in interpretability or merely artifacts of model-specific text processing.
+Until then, our results should be interpreted as proof-of-concept that linguistic editing can reduce measured disagreement, rather than as direct policy recommendations for communication strategy.